### 13 Feb 2024 v1

Abstract
We study the active few-shot fine-tuning of large neural networks to downstream tasks. We show that few-shot fine-tuning is an instance of a generalization of classical active learning, transductive active learning, and we propose ITL, short for information-based transductive learning, an approach which samples adaptively to maximize the information gained about specified downstream tasks. Under general regularity assumptions, we prove that ITL converges uniformly to the smallest possible uncertainty obtainable from the accessible data. To the best of our knowledge, we are the first to derive generalization bounds of this kind, and they may be of independent interest for active learning. We apply ITL to the few-shot fine-tuning of large neural networks and show that ITL substantially improves upon the state-of-the-art.
1 Introduction
Despite the remarkable successes of large neural networks (NNs) across various fields, such as image classification and natural language processing, their performance can deteriorate when faced with slight variations between the source and target domains (Recht et al., 2019; Hendrycks & Dietterich, 2019; Koh et al., 2021; Lee et al., 2022). Additionally, training large NNs requires large amounts of (labeled) data, which is often expensive or even impossible to obtain, and furthermore, training on such large datasets requires prohibitive computational resources.
Fine-tuning a large pre-trained model on a (small) dataset from the target domain is a cost- and computation-effective approach to address the distribution shift between source and target domains. While previous work has studied the effectiveness of various training procedures for fine-tuning (Howard & Ruder, 2018; Kornblith et al., 2019; Shen et al., 2021; Lee et al., 2022; Silva-Rodríguez et al., 2023), the problem of obtaining a good dataset for fine-tuning has received less attention. Selecting such a small dataset is challenging, as it requires selecting the most relevant and diverse data from a large dataset based only on a few reference examples from the target domain.
In this work, we propose a generalization of classical active learning, transductive active learning. We show that the fine-tuning of large neural networks can be seen as transductive active learning, and propose ITL which significantly improves upon the state-of-the-art — enabling efficient few-shot fine-tuning of large neural networks.
Transductive active learning
We consider the problem of transductive active learning, where provided a stochastic process $\{f(\boldsymbol{x})\}_{\boldsymbol{x}\in\mathcal{X}}$ over a domain $\mathcal{X}$ and given two arbitrary subsets of the domain; a target space $\mathcal{A}\subseteq\mathcal{X}$, and a sample space $\mathcal{S}\subseteq\mathcal{X}$, we study the question:
How can we learn $f$ within $\mathcal{A}$ by actively sampling observations within $\mathcal{S}$?
Prior work on active learning commonly aims to learn $f$ globally, i.e., across the entire domain $\mathcal{X}$. However, in many real-world problems, (i) the domain $\mathcal{X}$ is so large that learning $f$ globally is hopeless or (ii) agents have limited information and cannot access the entire domain $\mathcal{X}$ (e.g., due to restricted access or to act safely). Thus, global learning is often not desirable or even possible. Instead, intelligent systems are typically required to act in a more directed manner and extrapolate beyond their limited information. This work studies the above two aspects of active learning, which have remained largely unaddressed by prior work. We provide a comprehensive overview of related work in Section 5.
Figure 1: Instances of transductive active learning where the target space $\mathcal{A}$ is shown in blue and the sample space $\mathcal{S}$ is shown in gray. The points denote plausible observations within $\mathcal{S}$ to “learn” $\mathcal{A}$. In (A), the target space contains “everything” within $\mathcal{S}$. In (B, C, D), one makes observations “directed” towards learning about a particular target.
The fine-tuning of large neural networks is an instance of transductive active learning, where the target space represents the test distribution $\mathcal{P}_{\!\!\mathcal{A}}$ over which we want to minimize risk, and the sample space represents the train distribution $\mathcal{P}_{\!\mathcal{S}}$ from which we can sample data points to fine-tune our model. Figure 1 visualizes some instances of transductive active learning. Whereas most prior work has focused on the instance ${\mathcal{X}=\mathcal{A}=\mathcal{S}}$, transductive active learning generalizes to other instances such as (A), (B), (C), and (D).
Contributions
  • We prove the uniform convergence of uncertainty for ITL over the target space $\mathcal{A}$ to the smallest attainable value, given samples from the sample space $\mathcal{S}$ (Theorems 3.1 and 3.2). To the best of our knowledge, we are the first to derive such generalization bounds with rates in an active learning setting.
  • We apply the transductive active learning framework to batch-wise active few-shot fine-tuning of large neural networks and show empirically that ITL substantially outperforms the state-of-the-art (Section 4).
2 Preliminaries
[…]
We study the “adaptive” setting, where in round $n$ the agent selects a point $\boldsymbol{x}_{n}\in\mathcal{S}$ and observes ${y_{n}=y_{\boldsymbol{x}_{n}}}$. The agent’s choice of $\boldsymbol{x}_{n}$ may depend on the prior observations $\mathcal{D}_{n-1}\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\{(\boldsymbol{x}_{i},y_{i})\}_{i<n}$.
Background on information theory
We briefly recap several important concepts from information theory, of which we provide formal definitions in Appendix A. The (differential) entropy $\mathrm{H}\left[\boldsymbol{f}\right]$ is a measure of uncertainty about $\boldsymbol{f}$ and the conditional entropy $\mathrm{H}\!\left[\boldsymbol{f}\ \middle|\ \boldsymbol{y}\right]$ is the (expected) posterior uncertainty about $\boldsymbol{f}$ after observing $\boldsymbol{y}$. The information gain ${\mathrm{I}\left(\boldsymbol{f};\boldsymbol{y}\right)=\mathrm{H}\left[\boldsymbol{f}\right]-\mathrm{H}\!\left[\boldsymbol{f}\ \middle|\ \boldsymbol{y}\right]}$ measures the (expected) reduction in uncertainty about $\boldsymbol{f}$ due to $\boldsymbol{y}$. The maximum information gain from $n$ noisy observations within $\mathcal{S}$ is
$\gamma_{n}\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\max_{\begin{subarray}{c}X\subseteq\mathcal{S}\\ |X|\leq n\end{subarray}}\mathrm{I}\left(\boldsymbol{f}_{\!\!X};\boldsymbol{y}_{\!X}\right).$
This has been used previously (e.g., by Srinivas et al., 2009; Chowdhury & Gopalan, 2017; Vakili et al., 2021) as a measure of the “information capacity” of $f$.
3 Information-based Transductive Learning
We propose ITL, which greedily maximizes the information gain at each round $n$ between the prediction targets $\boldsymbol{f}_{\!\!\mathcal{A}}$ and the observation $y_{\boldsymbol{x}}$ conditioned on the prior observations $\mathcal{D}_{n-1}$. Formally,
$\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{S}}\mathrm{I}(\boldsymbol{f}_{\!\!\mathcal{A}};y_{\boldsymbol{x}}\mid\mathcal{D}_{n-1}).$ (ITL)
This simple decision rule generalizes several widely used algorithms which we discuss in more detail in Section 5. We also study the correlation-based decision rule,
$\boldsymbol{x}_{n}$ $=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{S}}\sum_{\boldsymbol{x^{\prime}}\in\mathcal{A}}\mathrm{Cor}{}\left[f_{\boldsymbol{x}},f_{\boldsymbol{x^{\prime}}}\mid\mathcal{D}_{n-1}\right]$ (CTL)
which we show to be a loose approximation of ITL in Appendix D. Unlike CTL, ITL takes into account the mutual dependence between points in $\mathcal{A}$.
In the following Sections 3.1 and 3.2, we discuss general settings in which the convergence properties of ITL can be analyzed theoretically.
3.1 Gaussian Process Setting
When $f\sim\mathcal{GP}({}\mu,k)$ is a Gaussian process (GP, Williams & Rasmussen (2006)) with known mean function $\mu$ and kernel $k$, and the noise $\varepsilon_{\boldsymbol{x}}$ is mutually independent and zero-mean Gaussian with known variance $\rho^{2}(\boldsymbol{x})>0$, the ITL objective has a closed form expression:
$\mathrm{I}(\boldsymbol{f}_{\!\!\mathcal{A}};y_{\boldsymbol{x}}\mid\mathcal{D}_{n-1})=\frac{1}{2}\log\left(\frac{\mathrm{Var}{}\left[y_{\boldsymbol{x}}\mid\mathcal{D}_{n-1}\right]missing}{\mathrm{Var}{}\left[y_{\boldsymbol{x}}\mid\boldsymbol{f}_{\!\!\mathcal{A}},\mathcal{D}_{n-1}\right]missing}\right).$ (1)
Further, the information capacity $\gamma_{n}$ is sublinear in $n$ for a rich class of GPs, with rates summarized in Table 3 of the appendix (Srinivas et al., 2009; Vakili et al., 2021). If ${|\mathcal{A}|<\infty}$ then ${\gamma_{n}(\mathcal{A};\mathcal{S})\leq\mathcal{O}\left(|\mathcal{A}|\log n\right)}$ is a loose upper bound. [^3] This follows from a simple reduction to an $|\mathcal{A}|$-dimensional linear bandit.
Convergence to irreducible uncertainty
So far, our discussion was centered around the role of the target space $\mathcal{A}$ in facilitating directed learning. An orthogonal contribution of this work is to study extrapolation from the sample space $\mathcal{S}$ to points ${\boldsymbol{x}\in\mathcal{A}\setminus\mathcal{S}}$. To this end, we derive bounds on the marginal posterior variance $\sigma_{n}^{2}(\boldsymbol{x})\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\mathrm{Var}{}\left[f(\boldsymbol{x})\mid\mathcal{D}_{n}\right]$ for points in $\mathcal{A}$. For ease of presentation, we assume in the following that $\mathcal{S}\subseteq\mathcal{A}$, however, the results are straightforward to generalize to other instances (Hübotter et al., 2024). For ITL, these bounds imply uniform convergence of the variance for a rich class of GPs. To the best of our knowledge, we are the first to derive such bounds, and they might be of independent interest for active learning.
We define the irreducible uncertainty as the variance of $f(\boldsymbol{x})$ provided complete knowledge of $f$ in $\mathcal{S}$:
$\eta^{2}(\boldsymbol{x};\mathcal{S})\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\mathrm{Var}{}[f_{\boldsymbol{x}}\mid\boldsymbol{f}_{\!\!\mathcal{S}}].$
As the name suggests, $\eta^{2}(\boldsymbol{x};\mathcal{S})$ represents the smallest uncertainty one can hope to achieve from observing only within $\mathcal{S}$. For all $\boldsymbol{x}\in\mathcal{S}$, it is easy to see that $\eta^{2}(\boldsymbol{x};\mathcal{S})=0$. However, the irreducible uncertainty of $\boldsymbol{x}\not\in\mathcal{S}$ may be (and typically is!) strictly positive.
Theorem 3.1 (Generalization bound on marginal variance for ITL).
Assume that ${f\sim\mathcal{GP}({}\mu,k)}$ with known mean function $\mu$ and kernel $k$, the noise $\varepsilon_{\boldsymbol{x}}$ is mutually independent and zero-mean Gaussian with known variance $\rho^{2}(\boldsymbol{x})>0$, and $\gamma_{n}$ is sublinear in $n$. Then, for any $n\geq 0,\epsilon>0$, and $\boldsymbol{x}\in\mathcal{A}$, there exists a constant $C_{\epsilon}$ independent of $n$ such that
[…]
where $\nu_{n,\epsilon}^{2}\operatorname*{\smash{\overset{\mathrm{def}}{=}}}C_{\epsilon}\gamma_{n}/n$ denotes the reducible uncertainty.
Intuitively, Equation 2 of Theorem 3.1 can be understood as bounding an epistemic “generalization gap” (Wainwright, 2019) of the learner. We provide a formal proof of Theorem 3.1 in Section C.3. By choosing, e.g., $\smash{\epsilon=c\sqrt{\gamma_{n}/n}}$ we have ${C_{\epsilon}\leq c^{\prime}\sqrt{n\gamma_{n}}}$ with constants $c$ and $c^{\prime}$ (cf. Section C.3.2), which implies that $\smash{\nu_{n,\epsilon}^{2}+\epsilon\leq c^{\prime}\sqrt{\gamma_{n}^{3}/n}}$. This guarantees that the reducible uncertainty converges, e.g., for Gaussian and smooth Matérn kernels.
3.2 Agnostic Setting
[…]
Theorem 3.2 (Bound on generalization error for ITL, following Abbasi-Yadkori (2013); Chowdhury & Gopalan (2017)).
Pick any $\delta\in(0,1)$ and $\epsilon>0$. Assume that $f^{\star}$ lies in the reproducing kernel Hilbert space $\mathcal{H}_{k}(\mathcal{X})$ of the kernel $k$ with norm $\smash{\left\|f^{\star}\right\|_{k}<\infty}$, the noise $\varepsilon_{n}$ is conditionally $\rho$-sub-Gaussian, and $\gamma_{n}$ is sublinear in $n$. Let $\smash{\beta_{n}(\delta)=\left\|f^{\star}\right\|_{k}+\rho\sqrt{2(\gamma_{n}+1+\log(1/\delta))}}$. Then, for any $n\geq 0$ and $\boldsymbol{x}\in\mathcal{A}$ jointly with probability at least $1-\delta$,
[…]
We provide a formal proof of Theorem 3.2 in Section C.4. Theorem 3.2 generalizes confidence bounds of prior works to the extrapolation setting, where some prediction targets $\boldsymbol{x}\in\mathcal{A}$ lie outside the sample space $\mathcal{S}$. Theorem 3.2 can therefore be interpreted as a generalization bound. For prediction targets ${\boldsymbol{x}\in\mathcal{A}\cap\mathcal{S}}$, the irreducible uncertainty vanishes, and we recover previous results.
4 Few-Shot Fine-Tuning of Neural Networks
As discussed in Section 1, the fine-tuning of large neural networks to downstream tasks can be framed as a transductive active learning problem. In our experiments, we consider a supervised learning problem, where the function $f$ maps inputs ${\boldsymbol{x}\in\mathcal{X}}$ to outputs ${y\in\mathcal{Y}}$ [^6] We remark that also the semi-supervised or unsupervised fine-tuning of models can be framed as a transductive active learning problem. We have access to noisy samples from a train data distribution $\mathcal{P}_{\!\mathcal{S}}$ on $\mathcal{X}$, and we would like to learn $f$ such that our estimate minimizes a given risk measure, such as classification error, with respect to a test distribution $\mathcal{P}_{\!\!\mathcal{A}}$ on $\mathcal{X}$. Ideally, $\mathcal{P}_{\!\mathcal{S}}$ and $\mathcal{P}_{\!\!\mathcal{A}}$ are identical, however, in many real-world cases $\mathcal{P}_{\!\mathcal{S}}$ differs substantially from $\mathcal{P}_{\!\!\mathcal{A}}$. This is for instance the case when samples from $\mathcal{P}_{\!\!\mathcal{A}}$ are expensive to obtain, such as in medical imaging. The goal is to actively and efficiently sample from $\mathcal{P}_{\!\mathcal{S}}$ to minimize risk with respect to $\mathcal{P}_{\!\!\mathcal{A}}$. In this setting, $\mathcal{P}_{\!\!\mathcal{A}}$ and $\mathcal{P}_{\!\mathcal{S}}$ are akin to $\mathcal{A}$ and $\mathcal{S}$, respectively. [^7] The setting with target distributions $\mathcal{P}_{\!\!\mathcal{A}}$ can be reduced to considering target sets $\mathcal{A}$, cf. Appendix E. We show in this section that ITL can efficiently learn $f$ using only few reference examples from $\mathcal{P}_{\!\!\mathcal{A}}$.
From neural networks to stochastic processes
Using a neural network to parameterize $f$, we can approximate $f$ by a stochastic process by extracting embeddings $\boldsymbol{\phi}(\boldsymbol{x})$, e.g., latent representations, from the network. Collecting them as rows in the design matrix $\boldsymbol{\Phi}$ of inputs $X$, we approximate the network by the linear function ${\boldsymbol{f}_{\!\!X}=\boldsymbol{\Phi}\boldsymbol{\theta}}$ with weights $\boldsymbol{\theta}$. Approximating the weights by ${\boldsymbol{\theta}\sim\mathcal{N}({}\boldsymbol{\mu},\boldsymbol{\Sigma})}$ implies $\smash{\boldsymbol{f}_{\!\!X}\sim\mathcal{N}({}\boldsymbol{\Phi}\boldsymbol{\mu},\boldsymbol{\Phi}\boldsymbol{\Sigma}\boldsymbol{\Phi}^{\top})}$. The embeddings $\boldsymbol{\phi}(\boldsymbol{x})$ are commonly taken either as latent representations encoded by $f$ (Lee et al., 2018; Devlin et al., 2019) or parameter gradients (Khan et al., 2019; He et al., 2020; Holzmüller et al., 2023) in which case $\smash{\boldsymbol{\Phi}\boldsymbol{\Phi}^{\top}}$ is the (empirical) neural tangent kernel (Arora et al., 2019). [^8] We extensively discuss embeddings in Section G.2. We use this Gaussian approximation with parameter gradient embeddings and ${\boldsymbol{\Sigma}=\boldsymbol{I}}$ to evaluate the ITL decision rule, and select inputs for labeling and training of $f$. The covariance matrix $\boldsymbol{\Sigma}$ quantifies uncertainty about the parameters $\boldsymbol{\theta}$, and we study approximations in Section G.3.
Batch selection via conditional embeddings
Efficient labeling and training necessitates a batch-wise selection of inputs. The selection of a batch of size $b>1$ can be seen as an individual non-adaptive active learning problem, and significant recent work has shown that batch diversity is crucial in this setting (Ash et al., 2020; Zanette et al., 2021; Holzmüller et al., 2023; Pacchiano et al., 2024). A batch-wise selection strategy is formalized by the following non-adaptive transductive active learning problem:
$B_{n}=\operatorname*{arg\,max}_{\begin{subarray}{c}B\subseteq\mathcal{S},|B|=b\end{subarray}}\mathrm{I}(\boldsymbol{f}_{\!\!\mathcal{A}};\boldsymbol{y}_{\!B}\mid\mathcal{D}_{n-1}).$ (4)
This problem is NP-hard (Krause & Golovin, 2014; Golovin & Krause, 2011), and ITL optimizes it greedily:
$\boldsymbol{x}_{n,i}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{S}}\mathrm{I}(\boldsymbol{f}_{\!\!\mathcal{A}};y_{\boldsymbol{x}}\mid\mathcal{D}_{n-1},\boldsymbol{y}_{\!\boldsymbol{x}_{n,1:i-1}}).$ (5)
We show in Appendix B that the approximation error of $B^{\prime}_{n}=\boldsymbol{x}_{n,1:b}$ can be bounded in terms of the submodularity ratio of ITL (Das & Kempe, 2018). For example, if $\mathcal{S}\subseteq\mathcal{A}$, the problem from Equation 4 is submodular and the $B^{\prime}_{n}$ is a $(1-1/e)$-approximation of $B_{n}$ (Nemhauser et al., 1978; Krause & Golovin, 2014). The batch $B_{n}$, and therefore also $B^{\prime}_{n}$, is diverse and informative by design. Prior work has shown that when the problem from Equation 4 is submodular, the greedy solution $B^{\prime}_{n}$ is also competitive with a fully sequential “batchless” decision rule (Chen & Krause, 2013; Esfandiari et al., 2021). We discuss an efficient iterative implementation of Equation 5 in Section G.4.
4.1 Experiments
As testbeds, we evaluate the performance of ITL on two image classification tasks: MNIST (LeCun et al., 1998) and CIFAR-100 (Krizhevsky et al., 2009). In both cases, we take $\mathcal{P}_{\!\mathcal{S}}$ to be the data distribution, and we consider the task of learning the digits $3$, $6$, and $9$ (MNIST) or the first $10$ categories of CIFAR-100 via few-shot training of NNs. [^9] That is, we restrict $\mathcal{P}_{\!\!\mathcal{A}}$ to the support of points with labels $\{3,6,9\}$ (MNIST) or labels $\{0,\dots,9\}$ (CIFAR-100) and train a neural network using few examples drawn from $\mathcal{P}_{\!\mathcal{S}}$.
Architectures
For MNIST, we train a simple convolutional neural network with ReLU activations, three convolutional layers with max-pooling, and two fully-connected layers. For CIFAR-100, we fine-tune an EfficientNet-B0 (Tan & Le, 2019) pre-trained on ImageNet (Deng et al., 2009), augmented by a final fully-connected layer. We train the NNs using the cross-entropy loss and the ADAM optimizer (Kingma & Ba, 2014).
Results
ITL and CTL consistently and significantly outperform Random, various widely used “undirected” heuristics such as BADGE (Ash et al., 2020), [^10] We show in Section G.5 that ITL and CTL also substantially outperform a wide selection of other “undirected” heuristics. and other “directed” heuristics such as CosineSimilarity and InformationDensity (Settles & Craven, 2008). In fact, the CosineSimilarity with $\mathcal{A}$ using embeddings $\boldsymbol{\phi}$ (cf. Section G.5), specializes CTL when ${\boldsymbol{\Sigma}=\boldsymbol{I}}$ and the batch size is $1$. Our main results are summarized in Figure 2.
[…]
[]nns
Figure 2: Few-shot training of NNs on MNIST (left) and CIFAR-100 (right). Random selects each observation uniformly at random from $\mathcal{P}_{\!\mathcal{S}}$. The batch size is $1$ for MNIST and $10$ for CIFAR-100. Uncertainty bands correspond to one standard error over $10$ random seeds. We see that ITL significantly outperforms the state-of-the-art, and in particular, retrieves substantially more samples from the support of $\mathcal{P}_{\!\!\mathcal{A}}$ than competing methods. This trend becomes even more pronounced in more difficult large-scale learning tasks (cf. Figure 7 in Appendix G). See Appendix G for details and additional experiments.
[…]
[]nns_batch_selection
Figure 3: Batch selection via conditional embeddings improves substantially over selecting the top-$b$ candidates proposed by the decision rule. This is the CIFAR-100 experiment (where $b=10$).
In Figure 3, we compare batch selection via conditional embeddings (BaCE) to selecting the top-$b$ points according to the decision rule (which does not yield diverse batches). We observe a significant improvement in accuracy and data retrieval when using BaCE. We expect the gap between both approaches to widen further with larger batch sizes.
Improved data retrieval
Both ITL and CTL use the learned embeddings to observe points similar to $\mathcal{P}_{\!\!\mathcal{A}}$, and in particular, obtain more samples from the support of $\mathcal{P}_{\!\!\mathcal{A}}$ than Random sampling of $\mathcal{P}_{\!\mathcal{S}}$. We observe empirically that ITL obtains samples from $\mathcal{P}_{\!\!\mathcal{A}}$ at more than twice the rate of CosineSimilarity, which translates to a significant improvement in accuracy in more difficult learning tasks, while requiring fewer (labeled) samples from $\mathcal{P}_{\!\mathcal{S}}$. This phenomenon manifests for both MNIST and CIFAR-100, as well as imbalanced data distributions $\mathcal{P}_{\!\mathcal{S}}$ or imbalanced reference samples from $\mathcal{P}_{\!\!\mathcal{A}}$ (cf. Section G.6). The improvement in accuracy appears to further increase in the large-data regime, where the learning tasks become more difficult (cf. Figure 7 in Appendix G).
5 Related Work
Data retrieval
Substantial prior work has studied data retrieval, e.g., in vector databases, using approximate nearest neighbor search (Johnson et al., 2019; Guo et al., 2020; Aumüller et al., 2020) where cosine similarity is a standard metric. Following the link between cosine similarity and CTL, ITL can be seen as a generalization of cosine similarity-based retrieval to batch and query sizes larger than one. Few-shot fine-tuning can be implemented also for sequences of different tasks (Vinyals et al., 2016), and as such task-driven few-shot fine-tuning can be seen as a form of “memory recall” akin to associative memory (Hopfield, 1982).
Undirected active learning
Our work extends several previous works in active learning and experimental design (Chaloner & Verdinelli, 1995; Settles, 2009). Uncertainty sampling (Lewis & Catlett, 1994, US) is one of the most popular active learning methods. US selects points $\boldsymbol{x}$ with high prior uncertainty: ${\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{S}}\sigma_{n-1}^{2}(\boldsymbol{x})}$. This is in stark contrast to ITL which selects points $\boldsymbol{x}$ that minimize posterior uncertainty about $\mathcal{A}$. It can be seen that US is the special “undirected” (i.e., $\mathcal{S}\subseteq\mathcal{A}$) case of ITL when the observation noise is homoscedastic (cf. Section C.2).
Several works have previously found entropy-based decision rules to be useful for undirected active learning (Krause & Guestrin, 2007; Guo & Greiner, 2007; Krause et al., 2008) and semi-supervised learning (Grandvalet & Bengio, 2004). A variance-based variant of ITL has previously been proposed by Yu et al. (2006) in the special case of global active learning without proving theoretical guarantees. Shoham & Avron (2023) recently analyzed experimental designs for global one-shot learning in overparameterized models.
Substantial work on active learning has studied entropy-based criteria in parameter-space, most notably BALD (Houlsby et al., 2011), which selects actions ${\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{X}}\mathrm{I}\!\left(\boldsymbol{\theta};y_{\boldsymbol{x}}\ \middle|\ \mathcal{D}_{n-1}\right)}$, where $\boldsymbol{\theta}$ is the random parameter vector of a parametric model (e.g., via Bayesian deep learning). Such methods are inherently “undirected” in the sense that they do not facilitate learning on specific prediction targets. In contrast, ITL operates in output-space where it is straightforward to specify prediction targets, and which is computationally easier.
Directed active learning
More recently, “directed” active learning methods have been proposed for the problem of determining the optimum of an unknown function, also known as best-arm identification (Audibert et al., 2010) or pure exploration bandits (Bubeck et al., 2009). Entropy search methods (Hennig & Schuler, 2012; Hernández-Lobato et al., 2014) are widely used and select ${\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{X}}\mathrm{I}\!\left(\boldsymbol{x}^{*};y_{\boldsymbol{x}}\ \middle|\ \mathcal{D}_{n-1}\right)}$ in input-space where ${\boldsymbol{x}^{*}=\operatorname*{arg\,max}_{\boldsymbol{x}}f_{\boldsymbol{x}}}$. Similarly to ITL, output-space entropy search methods (Hoffman & Ghahramani, 2015; Wang & Jegelka, 2017), which select ${\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{X}}\mathrm{I}\!\left(f^{*};y_{\boldsymbol{x}}\ \middle|\ \mathcal{D}_{n-1}\right)}$ with ${f^{*}=\max_{\boldsymbol{x}}f_{\boldsymbol{x}}}$, are more computationally tractable. In fact, entropy search methods are a special case of transductive active learning with target space $\mathcal{A}=\{\boldsymbol{x}^{*}\}$. When subsampling target spaces via Thompson sampling (Russo et al., 2018), ITL approximates these output-space methods. Bogunovic et al. (2016) analyze TruVar in the context of Bayesian optimization and level set estimation. TruVar is akin to a variance-based variant of ITL with a similar notion of “target space”, but their algorithm and analysis rely on a threshold scheme which requires that ${\mathcal{A}\subseteq\mathcal{S}}$. Fiez et al. (2019) introduce the transductive linear bandit problem, which is a special case of transductive active learning limited to a linear function class and with the objective of determining the maximum within an initial candidate set.
Loosely related works
The general principle of non-active “transductive learning” was introduced by Vapnik (2006). The notion of “target” from transductive active learning is akin to the notion of “task” in curriculum learning (Bengio et al., 2009; Graves et al., 2017; Soviany et al., 2022). The study of settings where the irreducible uncertainty is zero is related to the study of estimability in experimental design (Graybill, 1961; Mutny & Krause, 2022). In feature selection, selecting features that maximize information gain with respect to a to-be-predicted label is a standard approach (Peng et al., 2005; Vergara & Estévez, 2014; Beraha et al., 2019) which is akin to ITL. Transductive active learning is complimentary to other learning methodologies, such as semi-supervised learning (Gao et al., 2020), self-supervised learning (Shwartz-Ziv & LeCun, 2023; Balestriero et al., 2023), and meta-learning (Kaddour et al., 2020; Rothfuss et al., 2023).
6 Limitations
Better models
In this work, we focus solely on sequential decision-making given some model, rather than asking how one should construct such a model so that it is representative of the ground truth. Learning stochastic models and approximations thereof have been the subject of much recent work (Blundell et al., 2015; Maddox et al., 2019; Daxberger et al., 2021; Lin et al., 2023). Specifically, ITL benefits if the model $f$ captures the right “correlations” between points and if its “uncertainty” about the prediction at a specific point is accurate. [^11] For example, the results from Section 4 rely on embeddings being trained to capture the distribution shift.
Efficient computation of ITL with large $\mathcal{A}$
Naïve computation of the ITL decision rule for $m$ points within ${\mathcal{S}}$ takes $\mathcal{O}\left(|\mathcal{A}|^{3}+m\right)$ time and $\mathcal{O}\left(|\mathcal{A}|^{2}\right)$ space (cf. Appendix F) which is prohibitive for large $|\mathcal{A}|$. A practical solution for large $\mathcal{A}$ is to subsample $\mathcal{A}$ (such as in Section 4) within each round, and we provide a formal analysis of this approach for large $\mathcal{A}$ in Appendix E. [^12] In our NN experiments, we subsample $\mathcal{A}$ to a small set with size equal to the number of classes.
7 Conclusion and Future Work
We investigated the generalization of active learning to settings with concrete prediction targets and/or with limited information due to constrained sample spaces. We proved novel generalization bounds which may be of independent interest for active learning. Finally, we demonstrated empirically that ITL can substantially improve the efficiency of fine-tuning neural networks.
We believe that the transductive active learning methodology can also improve data-efficiency beyond supervised learning, such as in semi-supervised or self-supervised learning and reinforcement learning from human feedback (Dwaracherla et al., 2024), and leave this as an exciting direction for future work. Our demonstrated improvements in data retrieval suggest applicability beyond fine-tuning such as for in-context learning (Brown et al., 2020) and retrieval-augmented generation (Lewis et al., 2020) as well as pre-training.
Acknowledgements
This project was supported in part by the European Research Council (ERC) under the European Union’s Horizon 2020 research and Innovation Program Grant agreement no. 815943, the Swiss National Science Foundation under NCCR Automation, grant agreement 51NF40 180545, and by a grant of the Hasler foundation (grant nr: 21039).
We thank Armin Lederer, Johannes Kirschner, Jonas Rothfuss, Lars Loch, Manish Prajapat, Nicolas Emmenegger, Parnian Kassraie, and Scott Sussex for feedback. We thank Anton Baumann for helpful discussions.
Appendices
A general principle of “transductive learning” was already formulated by the famous computer scientist Vladimir Vapnik in the 20th century. Vapnik proposes the following “imperative for a complex world”:
When solving a problem of interest, do not solve a more general problem as an intermediate step. Try to get the answer that you really need but not a more general one.
– Vapnik (2006)
These appendices provide additional background, proofs, experiment details, and ablation studies.
Contents
[…]
0[2]
Appendix A Background
A.1 Information Theory
Throughout this work, $\log$ denotes the natural logarithm. Given random vectors $\boldsymbol{x}$ and $\boldsymbol{y}$, we denote by
$\mathrm{H}\left[\boldsymbol{x}\right]$ =$\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\mathbb{E}_{p(\boldsymbol{x})}{}\left[-\log p(\boldsymbol{x})\right],$
$\mathrm{H}\!\left[\boldsymbol{x}\ \middle|\ \boldsymbol{y}\right]$ =$\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\mathbb{E}_{p(\boldsymbol{x},\boldsymbol{y})}{}\left[-\log p(\boldsymbol{x}\mid\boldsymbol{y})\right],\quad\text{and}$
$\mathrm{I}\left(\boldsymbol{x};\boldsymbol{y}\right)$ =$\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\mathrm{H}\left[\boldsymbol{x}\right]-\mathrm{H}\!\left[\boldsymbol{x}\ \middle|\ \boldsymbol{y}\right]$
the (differential) entropy, conditional entropy, and information gain, respectively (Cover, 1999). [^13] One has to be careful to ensure that $\mathrm{I}\left(\boldsymbol{x};\boldsymbol{y}\right)$ exists, i.e., $|\mathrm{I}\left(\boldsymbol{x};\boldsymbol{y}\right)|<\infty$. We will assume that this is the case throughout this work. When $\boldsymbol{x}$ and $\boldsymbol{y}$ are jointly Gaussian, this is satisfied when the noise variance $\rho^{2}$ is positive.
A.2 Gaussian Processes
The stochastic process $f$ is a Gaussian process (GP, Williams & Rasmussen (2006)), denoted ${f\sim\mathcal{GP}({}\mu,k)}$, with mean function $\mu$ and kernel $k$ if for any finite subset ${X=\{\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{n}\}\subseteq\mathcal{X}}$, $\boldsymbol{f}_{\!\!X}\sim\mathcal{N}({}\boldsymbol{\mu}_{\!X},\boldsymbol{K}_{\!XX})$ is jointly Gaussian with mean vector $\boldsymbol{\mu}_{\!X}(i)=\mu(\boldsymbol{x}_{i})$ and covariance matrix $\boldsymbol{K}_{\!XX}(i,j)=k(\boldsymbol{x}_{i},\boldsymbol{x}_{j})$.
In the following, we formalize the assumptions from the GP setting (cf. Section 3.1).
Assumption A.1 (Gaussian prior).
We assume that ${f\sim\mathcal{GP}({}\mu,k)}$ with known mean function $\mu$ and kernel $k$.
Assumption A.2 (Gaussian noise).
We assume that the noise $\varepsilon_{\boldsymbol{x}}$ is mutually independent and zero-mean Gaussian with known variance $\rho^{2}(\boldsymbol{x})>0$. We write ${\boldsymbol{P}_{\!X}=\mathrm{diag}{}\left\{\rho^{2}(\boldsymbol{x}_{1}),\dots,\rho^{2}(\boldsymbol{x}_{n})\right\}}$.
Under assumptions A.1 and A.2, the posterior distribution of $f$ after observing points $X$ is $\mathcal{GP}({}\mu_{n},k_{n})$ with
$\mu_{n}(\boldsymbol{x})$ $=\mu(\boldsymbol{x})+\boldsymbol{K}_{\!\boldsymbol{x}X}(\boldsymbol{K}_{\!XX}+\boldsymbol{P}_{\!X})^{-1}(\boldsymbol{y}_{\!X}-\boldsymbol{\mu}_{\!X}),$
$k_{n}(\boldsymbol{x},\boldsymbol{x^{\prime}})$ $=k(\boldsymbol{x},\boldsymbol{x^{\prime}})-\boldsymbol{K}_{\!\boldsymbol{x}X}(\boldsymbol{K}_{\!XX}+\boldsymbol{P}_{\!X})^{-1}\boldsymbol{K}_{\!X\boldsymbol{x^{\prime}}},$
$\sigma_{n}^{2}(\boldsymbol{x})$ $=k_{n}(\boldsymbol{x},\boldsymbol{x}).$
For Gaussian random vectors $\boldsymbol{f}$ and $\boldsymbol{y}$, entropy is ${\mathrm{H}\left[\boldsymbol{f}\right]=\frac{n}{2}\log(2\pi e)+\frac{1}{2}\log\left|\mathrm{Var}{}\left[\boldsymbol{f}\right]\right|}$, information gain is ${\mathrm{I}\left(\boldsymbol{f};\boldsymbol{y}\right)=\frac{1}{2}\left(\log\left|\mathrm{Var}{}\left[\boldsymbol{y}\right]\right|-\log\left|\mathrm{Var}{}\left[\boldsymbol{y}\mid\boldsymbol{f}\right]\right|\right)}$ and
$\gamma_{n}$ $=\max_{\begin{subarray}{c}X\subseteq\mathcal{X}\\ |X|\leq n\end{subarray}}\frac{1}{2}\log\left|\boldsymbol{I}+\boldsymbol{P}_{\!X}^{-1}\boldsymbol{K}_{\!XX}\right|.$
Appendix B Provably Efficient Non-adaptive Batch Selection and its Relation to Submodularity
Recall the non-adaptive optimization problem
$B_{n,k}=\operatorname*{arg\,max}_{\begin{subarray}{c}B\subseteq\mathcal{S}\\ |B|=k\end{subarray}}\mathrm{I}\!\left(\boldsymbol{f}_{\!\!\mathcal{A}};\boldsymbol{y}_{\!B}\ \middle|\ \mathcal{D}_{n-1}\right)$
from Equation 4, and denote by ${B^{\prime}_{n,k}=\boldsymbol{x}_{n,1:k}}$ the greedy approximation from Equation 5. Note that the objective function
$F_{n}(B)\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\mathrm{I}\!\left(\boldsymbol{f}_{\!\!\mathcal{A}};\boldsymbol{y}_{\!B}\ \middle|\ \mathcal{D}_{n-1}\right)=\mathrm{H}\!\left[\boldsymbol{f}_{\!\!\mathcal{A}}\ \middle|\ \mathcal{D}_{n-1}\right]-\mathrm{H}\!\left[\boldsymbol{f}_{\!\!\mathcal{A}}\ \middle|\ \mathcal{D}_{n-1},\boldsymbol{y}_{\!B}\right]$
is non-negative and monotone (i.e., $F_{n}(B^{\prime})\leq F_{n}(B)$ for any ${B^{\prime}\subseteq B\subseteq\mathcal{S}}$) since conditional entropy is monotone.
Let $\Delta_{n}(\boldsymbol{x}\mid B)\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\Delta_{n}(\{\boldsymbol{x}\}\mid B)\operatorname*{\smash{\overset{\mathrm{def}}{=}}}F_{n}(B\cup\{\boldsymbol{x}\})-F(B)$ denote the marginal gain of $\boldsymbol{x}\in\mathcal{S}$ given $B\subseteq\mathcal{S}$ which simplifies to
$\Delta_{n}(\boldsymbol{x}\mid B)$ $=\mathrm{I}\!\left(\boldsymbol{f}_{\!\!\mathcal{A}};\boldsymbol{y}_{\!B},y_{\boldsymbol{x}}\ \middle|\ \mathcal{D}_{n-1}\right)-\mathrm{I}\!\left(\boldsymbol{f}_{\!\!\mathcal{A}};\boldsymbol{y}_{\!B}\ \middle|\ \mathcal{D}_{n-1}\right)$
$=\mathrm{H}\left[\boldsymbol{f}_{\!\!\mathcal{A}}\mid\mathcal{D}_{n-1},\boldsymbol{y}_{\!B}\right]-\mathrm{H}\left[\boldsymbol{f}_{\!\!\mathcal{A}}\mid\mathcal{D}_{n-1},\boldsymbol{y}_{\!B},y_{\boldsymbol{x}}\right]$
$=\mathrm{I}\left(\boldsymbol{f}_{\!\!\mathcal{A}};y_{\boldsymbol{x}}\mid\mathcal{D}_{n-1},\boldsymbol{y}_{\!B}\right),$
and which is precisely the objective function of ITL from Equation 5.
Batch selection via conditional embeddings approximates $B_{n,k}$
Building upon the theory of maximizing monotone submodular functions (Nemhauser et al., 1978; Krause & Golovin, 2014), Das & Kempe (2018) study greedy maximization under “approximate” submodularity:
Definition B.1 (Submodularity ratio of ITL).
Das & Kempe (2018) define the submodularity ratio of $F_{n}$ up to cardinality $k\geq 1$ as
$\kappa_{n}(k)\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\min_{\begin{subarray}{c}B\subseteq B^{\prime}_{n,k}\\ X\subseteq\mathcal{S}:|X|\leq k\\ B\cap X=\emptyset\end{subarray}}\frac{\sum_{\boldsymbol{x}\in X}\Delta_{n}(\boldsymbol{x}\mid B)}{\Delta_{n}(X\mid B)},$ (9)
where they define $\frac{0}{0}\equiv 1$.
As a special case of theorem 6 from Das & Kempe (2018), applying that $F_{n}$ is non-negative and monotone, we obtain the following result.
Corollary B.2 (Efficiency of batch selection via conditional embeddings).
For any $n,k\geq 1$, the greedy solution $B^{\prime}_{n,k}$ provides a $\left(1-e^{-\kappa_{n}(b)}\right)$-approximation of $B_{n,k}$.
The special “undirected” case
If $\mathcal{S}\subseteq\mathcal{A}$, it is well known (Srinivas et al., 2009) that $F_{n}$ is submodular, which implies that $\kappa_{n}(k)\geq 1$ for all $k\geq 1$.
Appendix C Proofs
C.1 Definitions
We write
  • $\sigma^{2}\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\max_{\boldsymbol{x}\in\mathcal{X}}\sigma_{0}^{2}(\boldsymbol{x})$, and
  • $\tilde{\sigma}^{2}\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\max_{\boldsymbol{x}\in\mathcal{X}}\sigma_{0}^{2}(\boldsymbol{x})+\rho^{2}(\boldsymbol{x})$.
C.2 Undirected Case of ITL
We briefly examine the important special case of ITL where $\mathcal{S}\subseteq\mathcal{A}$. In this setting, for all $\boldsymbol{x}\in\mathcal{S}$, the decision rule of ITL simplifies to
$\mathrm{I}(\boldsymbol{f}_{\!\!\mathcal{A}};y_{\boldsymbol{x}}\mid\mathcal{D}_{n})$ ${=}\mathrm{I}(\boldsymbol{f}_{\!\!\mathcal{A}\setminus\{\boldsymbol{x}\}};y_{\boldsymbol{x}}\mid f_{\boldsymbol{x}},\mathcal{D}_{n})+\mathrm{I}\!\left(f_{\boldsymbol{x}};y_{\boldsymbol{x}}\ \middle|\ \mathcal{D}_{n}\right)$
${=}\mathrm{I}(f_{\boldsymbol{x}};y_{\boldsymbol{x}}\mid\mathcal{D}_{n})$
$=\mathrm{H}[y_{\boldsymbol{x}}\mid\mathcal{D}_{n}]-\mathrm{H}[\varepsilon_{\boldsymbol{x}}]$
where $(i)$ follows from the chain rule of information gain and $\boldsymbol{x}\in\mathcal{S}\subseteq\mathcal{A}$; and $(ii)$ follows from the conditional independence $\boldsymbol{f}_{\!\!\mathcal{A}}\perp y_{\boldsymbol{x}}\mid f_{\boldsymbol{x}}$.
If additionally $f$ is a GP and the observation noise is homoscedastic, then
$\mathrm{H}[y_{\boldsymbol{x}}\mid\mathcal{D}_{n}]-\mathrm{H}\left[\varepsilon_{\boldsymbol{x}}\right]=\frac{1}{2}\log\left(1+\frac{\mathrm{Var}{}\left[f_{\boldsymbol{x}}\mid\mathcal{D}_{n}\right]missing}{\mathrm{Var}{}\left[\varepsilon_{\boldsymbol{x}}\right]missing}\right).$
Therefore, when $\mathcal{S}\subseteq\mathcal{A}$ and observation noise is homoscedastic, ITL is equivalent to uncertainty sampling.
C.3 Proof of Theorem 3.1
We will now prove Theorem 3.1 where we assume throughout that $\mathcal{S}\subseteq\mathcal{A}$. Let
$\Gamma_{n}\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\max_{\boldsymbol{x}\in\mathcal{S}}\mathrm{I}\!\left(\boldsymbol{f}_{\!\!\mathcal{A}};y_{\boldsymbol{x}}\ \middle|\ \mathcal{D}_{n}\right)$ (10)
denote the objective value of ITL during round $n+1$.
Before proving the convergence outside $\mathcal{S}$ in Section C.3.1, we first prove the convergence of the “step-wise uncertainty reduction” $\Gamma_{n}$.
Theorem C.1 (Bound of uncertainty reduction for ITL).
For any $n\geq 1$, if ITL generated the sequence $\{\boldsymbol{x}_{i}\}_{i=1}^{n}$ then
$\Gamma_{n-1}\leq\frac{\gamma_{n}}{n}.$ (11)
Proof.
We have
$\Gamma_{n-1}$ $=\frac{1}{n}\sum_{i=0}^{n-1}\Gamma_{n-1}$
${\leq}\frac{1}{n}\sum_{i=0}^{n-1}\Gamma_{i}$
${=}\frac{1}{n}\sum_{i=0}^{n-1}\mathrm{I}(\boldsymbol{f}_{\!\!\mathcal{A}};y_{\boldsymbol{x}_{i+1}}\mid\mathcal{D}_{i})$
${=}\frac{1}{n}\mathrm{I}(\boldsymbol{f}_{\!\!\mathcal{A}};\boldsymbol{y}_{\!\boldsymbol{x}_{1:n}})$
$\leq\frac{1}{n}\max_{\begin{subarray}{c}X\subseteq\mathcal{S}\\ |X|=n\end{subarray}}\mathrm{I}(\boldsymbol{f}_{\!\!\mathcal{A}};\boldsymbol{y}_{\!X})$
${\leq}\frac{1}{n}\max_{\begin{subarray}{c}X\subseteq\mathcal{S}\\ |X|=n\end{subarray}}\mathrm{I}(\boldsymbol{f}_{\!\!X};\boldsymbol{y}_{\!X})$
$=\frac{\gamma_{n}}{n}$
where $(i)$ follows from the monotonicity of conditional information gain since $\mathcal{S}\subseteq\mathcal{A}$ which implies $\mathcal{D}_{i}\perp y_{\boldsymbol{x}}\mid\boldsymbol{f}_{\!\!\mathcal{A}}$ for all $\boldsymbol{x}\in\mathcal{S}$; $(ii)$ uses the objective of ITL; $(iii)$ uses the chain rule of information gain; $(iv)$ uses ${\boldsymbol{y}_{\!X}\perp\boldsymbol{f}_{\!\!\mathcal{A}}\mid\boldsymbol{f}_{\!\!X}}$ and the data processing inequality. The conditional independences follow from the assumption that the observation noise is independent. ∎
We prove a simple consequence of Theorem C.1 for points $\boldsymbol{x}$ that are both in the target space $\mathcal{A}$ and in the sample space $\mathcal{S}$.
Lemma C.2 (Uniform bound of marginal variance within $\mathcal{S}$).
For any $n\geq 0$ and $\boldsymbol{x}\in\mathcal{A}\cap\mathcal{S}$,
$\sigma_{n}^{2}(\boldsymbol{x})\leq 2\tilde{\sigma}^{2}\cdot\Gamma_{n}.$ (12)
Proof.
We have
$\sigma_{n}^{2}(\boldsymbol{x})$ $=\mathrm{Var}{}\left[f_{\boldsymbol{x}}\mid\mathcal{D}_{n}\right]-\underbrace{\mathrm{Var}{}[f_{\boldsymbol{x}}\mid f_{\boldsymbol{x}},\mathcal{D}_{n}]missing}_{0}$
[…]
$=\mathrm{Var}{}\left[y_{\boldsymbol{x}}\mid\mathcal{D}_{n}\right]-\mathrm{Var}{}[y_{\boldsymbol{x}}\mid f_{\boldsymbol{x}},\mathcal{D}_{n}]$
${\leq}\tilde{\sigma}^{2}\log\left(\frac{\mathrm{Var}{}\left[y_{\boldsymbol{x}}\mid\mathcal{D}_{n}\right]missing}{\mathrm{Var}{}[y_{\boldsymbol{x}}\mid f_{\boldsymbol{x}},\mathcal{D}_{n}]missing}\right)$
$=2\tilde{\sigma}^{2}\cdot\mathrm{I}(f_{\boldsymbol{x}};y_{\boldsymbol{x}}\mid\mathcal{D}_{n})$
${\leq}2\tilde{\sigma}^{2}\cdot\mathrm{I}(\boldsymbol{f}_{\!\!\mathcal{A}};y_{\boldsymbol{x}}\mid\mathcal{D}_{n})$
${\leq}2\tilde{\sigma}^{2}\cdot\max_{\boldsymbol{x^{\prime}}\in\mathcal{S}}\mathrm{I}(\boldsymbol{f}_{\!\!\mathcal{A}};y_{\boldsymbol{x^{\prime}}}\mid\mathcal{D}_{n})$
$=2\tilde{\sigma}^{2}\cdot\Gamma_{n}$
where $(i)$ follows from the noise assumption (cf. assumption A.2); $(ii)$ follows from Lemma C.13 and using monotonicity of variance; $(iii)$ follows from $\boldsymbol{x}\in\mathcal{A}$ and monotonicity of information gain; and $(iv)$ follows from ${\boldsymbol{x}\in\mathcal{S}}$. ∎
C.3.1 Convergence outside $\mathcal{S}$
We will now show convergence of marginal variance to the irreducible uncertainty for points outside the sample space.
Our proof roughly proceeds as follows: We construct an “approximate Markov boundary” of $\boldsymbol{x}$ in $\mathcal{S}$, and show (1) that the size of this Markov boundary is independent of $n$, and (2) that a small uncertainty reduction within the Markov boundary implies that the marginal variances at the Markov boundary and(!) $\boldsymbol{x}$ are small.
Definition C.3 (Approximate Markov boundary).
For any $\epsilon>0$, $n\geq 0$, and $\boldsymbol{x}\in\mathcal{X}$, we denote by $B_{n,\epsilon}(\boldsymbol{x})$ the smallest (multi-)subset of $\mathcal{S}$ such that
$\mathrm{Var}{}[f_{\boldsymbol{x}}\mid\mathcal{D}_{n},\boldsymbol{y}_{\!B_{n,\epsilon}(\boldsymbol{x})}]$ $\leq\eta^{2}(\boldsymbol{x};\mathcal{S})+\epsilon.$ (13)
We call $B_{n,\epsilon}(\boldsymbol{x})$ an $\epsilon$-approximate Markov boundary of $\boldsymbol{x}$ in $\mathcal{S}$.
Equation 13 is akin to the notion of the smallest Markov blanket in $\mathcal{S}$ of some $\boldsymbol{x}\in\mathcal{X}$ (called a Markov boundary) which is the smallest set $\mathcal{B}\subseteq\mathcal{S}$ such that $f_{\boldsymbol{x}}\perp\boldsymbol{f}_{\!\!\mathcal{S}}\mid\boldsymbol{f}_{\!\!\mathcal{B}}$.
Lemma C.4 (Existence of an approximate Markov boundary).
For any $\epsilon>0$, let $k$ be the smallest integer satisfying
$\frac{\gamma_{k}}{k}\leq\frac{\epsilon\lambda_{\min}^{2}(\mathrm{Var}{}\left[\boldsymbol{f}_{\!\!\mathcal{S}}\right])}{2|\mathcal{S}|^{2}\sigma^{4}\tilde{\sigma}^{2}}.$ (14)
Then, for any $n\geq 0$ and $\boldsymbol{x}\in\mathcal{X}$, there exists an $\epsilon$-approximate Markov boundary $B_{n,\epsilon}(\boldsymbol{x})$ of $\boldsymbol{x}$ in $\mathcal{S}$ with size at most $k$.
Lemma C.4 shows that for any $\epsilon>0$ there exists a universal constant $b_{\epsilon}$ such that
$|B_{n,\epsilon}(\boldsymbol{x})|\leq b_{\epsilon}\qquad\forall n\geq 0,\boldsymbol{x}\in\mathcal{X}.$ (15)
We defer the proof of Lemma C.4 to Section C.3.3 where we also provide an algorithm to compute $B_{n,\epsilon}(\boldsymbol{x})$.
Lemma C.5.
For any $\epsilon>0$, $n\geq 0$, and $\boldsymbol{x}\in\mathcal{X}$,
$\sigma_{n}^{2}(\boldsymbol{x})\leq 2\sigma^{2}\cdot\mathrm{I}(f_{\boldsymbol{x}};\boldsymbol{y}_{\!B_{n,\epsilon}(\boldsymbol{x})}\mid\mathcal{D}_{n})+\eta^{2}(\boldsymbol{x};\mathcal{S})+\epsilon$ (16)
where $B_{n,\epsilon}(\boldsymbol{x})$ is an $\epsilon$-approximate Markov boundary of $\boldsymbol{x}$ in $\mathcal{S}$.
Proof.
We have
$\sigma_{n}^{2}(\boldsymbol{x})$ =$\end{multlined}\mathrm{Var}{}\left[f_{\boldsymbol{x}}\mid\mathcal{D}_{n}\right]-\eta^{2}(\boldsymbol{x};\mathcal{S})+\eta^{2}(\boldsymbol{x};\mathcal{S})$
[…]
$=2\sigma^{2}\cdot\mathrm{I}(f_{\boldsymbol{x}};\boldsymbol{y}_{\!B_{n,\epsilon}(\boldsymbol{x})}\mid\mathcal{D}_{n})+\eta^{2}(\boldsymbol{x};\mathcal{S})+\epsilon$
where $(i)$ follows from the defining property of an $\epsilon$-approximate Markov boundary (cf. Equation 13); and $(ii)$ follows from Lemma C.13 and using monotonicity of variance. ∎
Lemma C.6.
For any $\epsilon>0,n\geq 0$, and $\boldsymbol{x}\in\mathcal{A}$,
$\mathrm{I}(f_{\boldsymbol{x}};\boldsymbol{y}_{\!B_{n,\epsilon}(\boldsymbol{x})}\mid\mathcal{D}_{n})\leq b_{\epsilon}\Gamma_{n}$ (17)
where $B_{n,\epsilon}(\boldsymbol{x})$ is an $\epsilon$-approximate Markov boundary of $\boldsymbol{x}$ in $\mathcal{S}$ with $|B_{n,\epsilon}(\boldsymbol{x})|\leq b_{\epsilon}$.
Proof.
We use the abbreviated notation ${B=B_{n,\epsilon}(\boldsymbol{x})}$. We have
$\mathrm{I}(f_{\boldsymbol{x}};\boldsymbol{y}_{\!B}\mid\mathcal{D}_{n})$ ${\leq}\mathrm{I}(\boldsymbol{f}_{\!\!\mathcal{A}};\boldsymbol{y}_{\!B}\mid\mathcal{D}_{n})$
${\leq}\sum_{\tilde{\boldsymbol{x}}\in B}\mathrm{I}(\boldsymbol{f}_{\!\!\mathcal{A}};y_{\tilde{\boldsymbol{x}}}\mid\mathcal{D}_{n})$
${\leq}b_{\epsilon}\max_{\tilde{\boldsymbol{x}}\in B}\mathrm{I}(\boldsymbol{f}_{\!\!\mathcal{A}};y_{\tilde{\boldsymbol{x}}}\mid\mathcal{D}_{n})$
${\leq}b_{\epsilon}\max_{\tilde{\boldsymbol{x}}\in\mathcal{S}}\mathrm{I}(\boldsymbol{f}_{\!\!\mathcal{A}};y_{\tilde{\boldsymbol{x}}}\mid\mathcal{D}_{n})$
$=b_{\epsilon}\Gamma_{n}$
where $(i)$ follows from monotonicity of mutual information; $(ii)$ follows from the chain rule of information gain and the monotonicity of conditional information gain which is due to $\mathcal{D}_{n}\perp y_{\tilde{\boldsymbol{x}}}\mid\boldsymbol{f}_{\!\!\mathcal{A}}$ for all $\tilde{\boldsymbol{x}}\in B\subseteq\mathcal{S}$; $(iii)$ follows from $b\leq b_{\epsilon}$; and $(iv)$ follows from $B\subseteq\mathcal{S}$. ∎
Proof of Theorem 3.1.
Fix any ${\boldsymbol{x}\in\mathcal{A}}$ and ${\epsilon>0}$. By Lemma C.4, there exists an $\epsilon$-approximate Markov boundary $B_{n,\epsilon}(\boldsymbol{x})$ of $\boldsymbol{x}$ in $\mathcal{S}$ such that $|B_{n,\epsilon}(\boldsymbol{x})|\leq b_{\epsilon}$. We have
$\sigma_{n}^{2}(\boldsymbol{x})$ ${\leq}2\sigma^{2}\cdot\mathrm{I}(f_{\boldsymbol{x}};\boldsymbol{y}_{\!B_{n,\epsilon}(\boldsymbol{x})}\mid\mathcal{D}_{n})+\eta^{2}(\boldsymbol{x};\mathcal{S})+\epsilon$
${\leq}2\sigma^{2}b_{\epsilon}\Gamma_{n}+\eta^{2}(\boldsymbol{x};\mathcal{S})+\epsilon$
${\leq}2\sigma^{2}b_{\epsilon}\frac{\gamma_{n+1}}{{n+1}}+\eta^{2}(\boldsymbol{x};\mathcal{S})+\epsilon$
where $(i)$ follows from Lemma C.5; $(ii)$ follows from Lemma C.6; and follows from Theorem C.1 using that ITL generated the sequence $\{\boldsymbol{x}_{i}\}_{i=1}^{n}$. The result follows by setting $C_{\epsilon}=2\sigma^{2}b_{\epsilon}$. ∎
C.3.2 Exemplary Application of Theorem 3.1
Assume that $\gamma_{n}$ is sublinear in $n$. Let $\epsilon=c\sqrt{\frac{\gamma_{n}}{n}}$ with ${c=2|\mathcal{S}|^{2}\sigma^{4}\tilde{\sigma}^{2}/\lambda_{\min}^{2}(\mathrm{Var}{}\left[\boldsymbol{f}_{\!\!\mathcal{S}}\right])}$. Then, by Equation 14, $b_{\epsilon}$ is the smallest integer satisfying
$\frac{\gamma_{b_{\epsilon}}}{b_{\epsilon}}\leq\sqrt{\frac{\gamma_{n}}{n}}.$
Note that $b_{\epsilon}$ depends on $n$, but can be bounded for instance by $\sqrt{n\gamma_{n}}$, since
$\frac{\gamma_{\sqrt{n\gamma_{n}}}}{\sqrt{n\gamma_{n}}}\leq\frac{\gamma_{\sqrt{n^{2}}}}{\sqrt{n\gamma_{n}}}\leq\frac{\gamma_{n}}{\sqrt{n\gamma_{n}}}=\sqrt{\frac{\gamma_{n}}{n}}$
where we use that $\gamma_{n}\leq n$ for sufficiently large $n$. Together with Theorem C.1 this implies for ITL that
$\nu_{n,\epsilon}^{2}+\epsilon$ $\leq 2\sigma^{2}\sqrt{n\gamma_{n}}\,\Gamma_{n}+c\sqrt{\gamma_{n}/n}$
$\leq c^{\prime}\sqrt{\gamma_{n}^{3}/n}$
for a constant $c^{\prime}$, e.g., $c^{\prime}=2\sigma^{2}+c$. This guarantees that the reducible uncertainty of ITL converges, e.g., for Gaussian and smooth Matérn kernels.
C.3.3 Existence of an Approximate Markov Boundary
We now derive Lemma C.4 which shows the existence of an approximate Markov boundary of $\boldsymbol{x}$ in $\mathcal{S}$.
Lemma C.7.
For any $k\geq 0$, there exists ${B\subseteq\mathcal{S}}$ with ${|B|=k}$ such that for all ${\boldsymbol{x^{\prime}}\in\mathcal{S}}$,
$\mathrm{Var}{}\left[f_{\boldsymbol{x^{\prime}}}\mid\boldsymbol{y}_{\!B}\right]\leq 2\tilde{\sigma}^{2}\frac{\gamma_{k}}{k}.$ (18)
Proof.
We choose $B\subseteq\mathcal{S}$ greedily using the acquisition function
$\tilde{\boldsymbol{x}}_{k}\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\operatorname*{arg\,max}_{\tilde{\boldsymbol{x}}\in\mathcal{S}}\mathrm{I}(\boldsymbol{f}_{\!\!\mathcal{S}};y_{\tilde{\boldsymbol{x}}}\mid\boldsymbol{y}_{\!B_{k-1}})$
where $B_{k}=\tilde{\boldsymbol{x}}_{1:k}$. Note that this is the special case of ITL with $\mathcal{S}\subseteq\mathcal{A}$, and hence, we have
$\mathrm{Var}{}\left[f_{\boldsymbol{x^{\prime}}}\mid\boldsymbol{y}_{\!B_{k}}\right]$ ${\leq}2\tilde{\sigma}^{2}\Gamma_{k}$
${\leq}2\tilde{\sigma}^{2}\frac{\gamma_{k+1}}{{k+1}}$
where $(i)$ is due to Lemma C.2; and $(ii)$ is due to Theorem C.1 and $\alpha_{k}(\mathcal{S};\mathcal{S})\leq 1$. ∎
Lemma C.8.
Given any $\epsilon>0$ and $B\subseteq\mathcal{S}$, such that for any $\boldsymbol{x^{\prime}}\in\mathcal{S}$,
$\mathrm{Var}{}\left[f_{\boldsymbol{x^{\prime}}}\mid\boldsymbol{y}_{\!B}\right]\leq\frac{\epsilon\lambda_{\min}^{2}(\mathrm{Var}{}\left[\boldsymbol{f}_{\!\!\mathcal{S}}\right])}{|\mathcal{S}|^{2}\sigma^{4}}.$ (19)
Then for any ${\boldsymbol{x}\in\mathcal{X}}$,
$\mathrm{Var}{}\left[f_{\boldsymbol{x}}\mid\boldsymbol{y}_{\!B}\right]\leq\mathrm{Var}{}\left[f_{\boldsymbol{x}}\mid\boldsymbol{f}_{\!\!\mathcal{S}}\right]+\epsilon.$ (20)
Proof.
We will denote the right-hand side of Equation 19 by $\epsilon^{\prime}$. We have
$\mathrm{Var}{}\left[f_{\boldsymbol{x}}\mid\boldsymbol{y}_{\!B}\right]$
[…]
${=}\mathrm{Var}_{f_{\boldsymbol{x}}}{}\left[f_{\boldsymbol{x}}\mid\boldsymbol{f}_{\!\!\mathcal{S}},\boldsymbol{y}_{\!B}\right]+\mathrm{Var}_{\boldsymbol{f}_{\!\!\mathcal{S}}}{}\left[\mathbb{E}_{f_{\boldsymbol{x}}}{}\left[f_{\boldsymbol{x}}\mid\boldsymbol{f}_{\!\!\mathcal{S}},\boldsymbol{y}_{\!B}\right]\mid\boldsymbol{y}_{\!B}\right]$
${=}\underbrace{\mathrm{Var}_{f_{\boldsymbol{x}}}{}\left[f_{\boldsymbol{x}}\mid\boldsymbol{f}_{\!\!\mathcal{S}}\right]missing}_{\text{irreducible uncertainty}}+\underbrace{\mathrm{Var}_{\boldsymbol{f}_{\!\!\mathcal{S}}}{}\left[\mathbb{E}_{f_{\boldsymbol{x}}}{}\left[f_{\boldsymbol{x}}\mid\boldsymbol{f}_{\!\!\mathcal{S}}\right]\mid\boldsymbol{y}_{\!B}\right]missing}_{\text{reducible (epistemic) uncertainty}}$
where $(i)$ follows from the law of total variance; $(ii)$ uses that the conditional variance of a Gaussian depends only on the location of observations and not on their value; and $(iii)$ follows from $f_{\boldsymbol{x}}\perp\boldsymbol{y}_{\!B}\mid\boldsymbol{f}_{\!\!\mathcal{S}}$ since $B\subseteq\mathcal{S}$. It remains to bound the reducible uncertainty.
Let $h:\mathbb{R}^{d}\to\mathbb{R},\;\boldsymbol{f}_{\!\!\mathcal{S}}\mapsto\mathbb{E}{}\left[f_{\boldsymbol{x}}\mid\boldsymbol{f}_{\!\!\mathcal{S}}\right]$ where we write $d\operatorname*{\smash{\overset{\mathrm{def}}{=}}}|\mathcal{S}|$. Using the formula for the GP posterior mean, we have
$h(\boldsymbol{f}_{\!\!\mathcal{S}})=\mathbb{E}{}\left[f_{\boldsymbol{x}}\right]+\boldsymbol{z}^{\top}(\boldsymbol{f}_{\!\!\mathcal{S}}-\mathbb{E}{}\left[\boldsymbol{f}_{\!\!\mathcal{S}}\right])$
where $\boldsymbol{z}\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\boldsymbol{K}^{-1}\boldsymbol{k}$, $\boldsymbol{K}\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\mathrm{Var}{}\left[\boldsymbol{f}_{\!\!\mathcal{S}}\right]$, and $\boldsymbol{k}\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\mathrm{Cov}{}\left[\boldsymbol{f}_{\!\!\mathcal{S}},f_{\boldsymbol{x}}\right]$. Because $h$ is a linear function in $\boldsymbol{f}_{\!\!\mathcal{S}}$ we have for the reducible uncertainty that
$\mathrm{Var}_{\boldsymbol{f}_{\!\!\mathcal{S}}}{}\left[h(\boldsymbol{f}_{\!\!\mathcal{S}})\mid\boldsymbol{y}_{\!B}\right]$ $=\boldsymbol{z}^{\top}\mathrm{Var}{}\left[\boldsymbol{f}_{\!\!\mathcal{S}}\mid\boldsymbol{y}_{\!B}\right]\boldsymbol{z}$
${\leq}d\cdot\boldsymbol{z}^{\top}\mathrm{diag}{}\left\{\mathrm{Var}{}\left[\boldsymbol{f}_{\!\!\mathcal{S}}\mid\boldsymbol{y}_{\!B}\right]\right\}\boldsymbol{z}$
${\leq}\epsilon^{\prime}d\left\|\boldsymbol{z}\right\|_{2}^{2}$
$=\epsilon^{\prime}d\left\|\boldsymbol{K}^{-1}\boldsymbol{k}\right\|_{2}^{2}$
$\leq\epsilon^{\prime}d\left\|\boldsymbol{K}^{-1}\right\|_{2}^{2}\left\|\boldsymbol{k}\right\|_{2}^{2}$
$=\frac{\epsilon^{\prime}d\left\|\boldsymbol{k}\right\|_{2}^{2}}{\lambda_{\min}^{2}(\boldsymbol{K})}$
where $(i)$ follows from Lemma C.12; and $(ii)$ follows from the assumption that $\mathrm{Var}{}\left[f_{\boldsymbol{x^{\prime}}}\mid\boldsymbol{y}_{\!B}\right]\leq\epsilon^{\prime}$ for all $\boldsymbol{x^{\prime}}\in\mathcal{S}$. We have
$\left\|\boldsymbol{k}\right\|_{2}^{2}$ $=\sum_{\begin{subarray}{c}\boldsymbol{x^{\prime}}\in\mathcal{S}\end{subarray}}\underbrace{\mathrm{Cov}{}\left[f_{\boldsymbol{x}},f_{\boldsymbol{x^{\prime}}}\right]^{2}}_{\leq\sigma^{4}}\leq d\sigma^{4}.$
Thus,
$\mathrm{Var}_{\boldsymbol{f}_{\!\!\mathcal{S}}}{}\left[h(\boldsymbol{f}_{\!\!\mathcal{S}})\mid\boldsymbol{y}_{\!B}\right]\leq\frac{\epsilon^{\prime}d^{2}\sigma^{4}}{\lambda_{\min}^{2}(\boldsymbol{K})}=\epsilon.$
Proof of Lemma C.4.
Let $B$ be the set of size $k$ generated by Lemma C.7 to satisfy ${\mathrm{Var}{}\left[f_{\boldsymbol{x^{\prime}}}\mid\boldsymbol{y}_{\!B}\right]\leq 2\tilde{\sigma}^{2}\gamma_{k}/k}$ for all $\boldsymbol{x^{\prime}}\in\mathcal{S}$. We have for any ${\boldsymbol{x}\in\mathcal{X}}$,
$\mathrm{Var}{}\left[f_{\boldsymbol{x}}\mid\mathcal{D}_{n},\boldsymbol{y}_{\!B}\right]$ ${\leq}\mathrm{Var}{}\left[f_{\boldsymbol{x}}\mid\boldsymbol{y}_{\!B}\right]$
${\leq}\mathrm{Var}{}\left[f_{\boldsymbol{x}}\mid\boldsymbol{f}_{\!\!\mathcal{S}}\right]+\epsilon$
where $(i)$ follows from monotonicity of variance; and $(ii)$ follows from Lemma C.8 and using the condition on $k$. ∎
We remark that Lemma C.7 provides an algorithm (just “undirected” ITL!) to compute an approximate Markov boundary, and the set $B$ returned by this algorithm is a valid approximate Markov boundary for all $\boldsymbol{x}\in\mathcal{X}$.
C.4 Proof of Theorem 3.2
We first formalize the assumptions of Theorem 3.2:
Assumption C.9 (Regularity of $f^{\star}$).
We assume that $f^{\star}$ is in a reproducing kernel Hilbert space $\mathcal{H}_{k}(\mathcal{X})$ associated with a kernel $k$ and has bounded norm, that is, $\left\|f\right\|_{k}\leq B$ for some finite $B\in\mathbb{R}$.
Assumption C.10 (Sub-Gaussian noise).
We further assume that each $\epsilon_{n}$ from the noise sequence $\{\varepsilon_{n}\}_{n=1}^{\infty}$ is conditionally zero-mean $\rho(\boldsymbol{x}_{n})$-sub-Gaussian with known constants ${\rho(\boldsymbol{x})>0}$ for all $\boldsymbol{x}\in\mathcal{X}$. Concretely,
$\forall n\geq 1,\lambda\in\mathbb{R}:\quad\mathbb{E}{}\!\left[e^{\lambda\epsilon_{n}}\ \middle|\ \mathcal{F}_{n-1}\right]\leq\exp\left(\frac{\lambda^{2}\rho^{2}(\boldsymbol{x}_{n})}{2}\right)$
where $\mathcal{F}_{n-1}$ denotes the $\sigma$-algebra generated by the random variables $\{\boldsymbol{x}_{i},\epsilon_{i}\}_{i=1}^{n-1}$ and $\boldsymbol{x}_{n}$.
We make use of the following foundational result, showing that under the above two assumptions the (misspecified) Gaussian process model from Section 3.1 is an all-time well-calibrated model of $f^{\star}$:
Lemma C.11 (Well-calibrated confidence intervals; Abbasi-Yadkori (2013); Chowdhury & Gopalan (2017)).
Pick ${\delta\in(0,1)}$ and let assumptions C.9 and C.10 hold. Let
$\beta_{n}(\delta)=\beta_{n}(\delta)=\left\|f^{\star}\right\|_{k}+\rho\sqrt{2(\gamma_{n}+1+\log(1/\delta))}$
[…]
$|f^{\star}(\boldsymbol{x})-\mu_{n}(\boldsymbol{x})|\leq\beta_{n}(\delta)\cdot\sigma_{n}(\boldsymbol{x})$
where $\mu_{n}(\boldsymbol{x})$ and $\sigma_{n}^{2}(\boldsymbol{x})$ are mean and variance (as defined in Section A.2) of the GP posterior of $f(\boldsymbol{x})$ conditional on the observations $\mathcal{D}_{n}$, pretending that $\varepsilon_{i}$ is Gaussian with variance $\rho^{2}(\boldsymbol{x}_{i})$.
The proof of Theorem 3.2 is a straightforward application of Lemmas C.11 and 3.1:
Proof of Theorem 3.2.
By Theorem 3.1, we have that for all ${\boldsymbol{x}\in\mathcal{A}}$,
$\sigma_{n}(\boldsymbol{x})$ $\leq\sqrt{\eta^{2}(\boldsymbol{x};\mathcal{S})+\nu_{n,\epsilon^{2}}^{2}+\epsilon^{2}}\leq\eta(\boldsymbol{x};\mathcal{S})+\nu_{n,\epsilon^{2}}+\epsilon.$
The result then follows by application of Lemma C.11. ∎
C.5 Useful Facts and Inequalities
We denote by $\preceq$ the Loewner partial ordering of symmetric matrices.
Lemma C.12.
Let $\boldsymbol{A}\in\mathbb{R}^{n\times n}$ be a positive definite matrix with diagonal $\boldsymbol{D}$. Then, $\boldsymbol{A}\preceq n\boldsymbol{D}$.
Proof.
Equivalently, one can show $n\boldsymbol{D}-\boldsymbol{A}\succeq\boldsymbol{0}$. We write $\boldsymbol{A}\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\boldsymbol{D}^{\nicefrac{{1}}{{2}}}\boldsymbol{Q}\boldsymbol{D}^{\nicefrac{{1}}{{2}}}$, and thus, $\boldsymbol{Q}=\boldsymbol{D}^{-\nicefrac{{1}}{{2}}}\boldsymbol{A}\boldsymbol{D}^{-\nicefrac{{1}}{{2}}}$ is a positive definite symmetric matrix with all diagonal elements equal to $1$. It remains to show that
$n\boldsymbol{D}-\boldsymbol{A}=\boldsymbol{D}^{\nicefrac{{1}}{{2}}}(n\boldsymbol{I}-\boldsymbol{Q})\boldsymbol{D}^{\nicefrac{{1}}{{2}}}\succeq\boldsymbol{0}.$
Note that $\sum_{i=1}^{n}\lambda_{i}(\boldsymbol{Q})=\mathrm{tr}\;\boldsymbol{Q}=n$, and hence, all eigenvalues of $\boldsymbol{Q}$ belong to $(0,n)$. ∎
Lemma C.13.
If $a,b\in(0,M]$ for some $M>0$ and $b\geq a$ then
$b-a\leq M\cdot\log\left(\frac{b}{a}\right).$ (21)
If additionally, $a\geq M^{\prime}$ for some $M^{\prime}>0$ then
$b-a\geq M^{\prime}\cdot\log\left(\frac{b}{a}\right).$ (22)
Proof.
Let $f(x)\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\log x$. By the mean value theorem, there exists $c\in(a,b)$ such that
$\frac{1}{c}=f^{\prime}(c)=\frac{f(b)-f(a)}{b-a}=\frac{\log b-\log a}{b-a}=\frac{\log(\frac{b}{a})}{b-a}.$
Thus,
$b-a=c\cdot\log\left(\frac{b}{a}\right)<M\cdot\log\left(\frac{b}{a}\right).$
Under the additional condition that $a\geq M^{\prime}$, we obtain
$b-a=c\cdot\log\left(\frac{b}{a}\right)>M^{\prime}\cdot\log\left(\frac{b}{a}\right).$
Appendix D Correlation-based Transductive Learning
We will briefly look at the decision rule
$\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{S}}\sum_{\boldsymbol{x^{\prime}}\in\mathcal{A}}\mathrm{Cor}{}\left[f_{\boldsymbol{x}},f_{\boldsymbol{x^{\prime}}}\mid\mathcal{D}_{n-1}\right]$ (23)
which as we will see can be thought of as loose approximation to ITL We refer to this decision rule as CTL, short for correlation-based transductive learning.
We consider the GP setting, i.e., assumptions A.1 and A.2 hold.
Approximation of ITL
The ITL objective can be shown to be lower bounded by
$\mathrm{I}\!\left(\boldsymbol{f}_{\!\!\mathcal{A}};y_{\boldsymbol{x}}\ \middle|\ \mathcal{D}_{n-1}\right)$ ${\geq}\frac{1}{|\mathcal{A}|}\sum_{\boldsymbol{x^{\prime}}\in\mathcal{A}}\mathrm{I}(f_{\boldsymbol{x^{\prime}}};y_{\boldsymbol{x}}\mid\mathcal{D}_{n-1})$
$=\frac{1}{2|\mathcal{A}|}\sum_{\boldsymbol{x^{\prime}}\in\mathcal{A}}\mathrm{I}\!\left(f_{\boldsymbol{x^{\prime}}};y_{\boldsymbol{x}}\ \middle|\ \mathcal{D}_{n-1}\right)$
${=}-\sum_{\boldsymbol{x^{\prime}}\in\mathcal{A}}\log\left(1-\mathrm{Cor}{}\left[f_{\boldsymbol{x^{\prime}}},y_{\boldsymbol{x}}\mid\mathcal{D}_{n-1}\right]^{2}\right)$
${\gtrsim}\sum_{\boldsymbol{x^{\prime}}\in\mathcal{A}}\mathrm{Cor}{}\left[f_{\boldsymbol{x^{\prime}}},y_{\boldsymbol{x}}\mid\mathcal{D}_{n-1}\right]^{2}+\mathrm{const}$
where $(i)$ is detailed in example 8.5.1 of (Cover, 1999); and $(iii)$ follows from Jensen’s inequality and exponentiating. $(i)$ is a loose approximation which highlights that ITL takes into account the correlations between points in $\mathcal{A}$ while CTL dooes not.
Appendix E Subsampling Target Spaces
When the target space $\mathcal{A}$ is large, it may be computationally infeasible to compute the exact objective. A natural approach to address this issue is to approximate the target space by a smaller set of size $K$.
One possibility is to select the $K$ “most explanatory” points within $\mathcal{A}$. This selection problem is similar to the batch selection problem in active learning (Holzmüller et al., 2023) and can be tackled using, e.g., the greedy max determinant or greedy max kernel distance strategies. In the remainder of this section, we study an alternative approach which is based on sampling points from $\mathcal{A}$ according to some probability distribution $\mathcal{P}_{\!\!\mathcal{A}}$ supported on $\mathcal{A}$.
E.1 Stochastic Target Spaces
Concretely, in iteration $n$, a subset $\mathcal{A}^{\prime}_{n}$ of $K$ points is sampled independently from $\mathcal{A}$ according to the distribution $\mathcal{P}_{\!\!\mathcal{A}}$ and the objective is computed on this subset.
It is straightforward to extend the convergence guarantees to the setting of stochastic target spaces. Intuitively, after sufficiently many iterations, each $\boldsymbol{x}\in\mathcal{A}$ will have been sampled roughly as often as its “weight” (i.e., probability) suggests.
Lemma E.1.
Consider the stochastic target space setting. For any $n\geq 0,K\geq 1$, and $\boldsymbol{x}\in\mathcal{A}$, with probability at least ${1-\exp(-n\nu/8)}$ where $\nu=1-(1-\mathcal{P}_{\!\!\mathcal{A}}(\boldsymbol{x}))^{K}$, we have that after at most $n$ iterations $\boldsymbol{x}$ was within the subsampled target space in at least $n\nu/2$ iterations.
Proof.
Let $Y_{i}\sim\mathrm{Binom}(K,\mathcal{P}_{\!\!\mathcal{A}}(\boldsymbol{x}))$ denote the random variable counting the number of occurrences of $\boldsymbol{x}$ in $\mathcal{A}^{\prime}_{i}$. Moreover, we write ${X_{i}\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\mathbbm{1}\{{\boldsymbol{x}\in\mathcal{A}^{\prime}_{i}}\}}$. Note that
$\nu_{i}\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\mathbb{E}{}X_{i}$ $=\mathbb{P}\left(\boldsymbol{x}\in\mathcal{A}^{\prime}_{i}\right)$
$=1-\mathbb{P}\left(Y_{i}=0\right)$
$=1-(1-\mathcal{P}_{\!\!\mathcal{A}}(\boldsymbol{x}))^{K}=\nu.$
Let $X\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\sum_{i=1}^{n}X_{i}$ with $\mathbb{E}{}X=n\nu$. By Chernoff’s bound,
$\mathbb{P}\left(X\leq\frac{n\nu}{2}\right)\leq\exp\left(-\frac{n\nu}{8}\right).$
With this result, the previously derived convergence guarantees follow also for the stochastic target space setting.
Appendix F Computational Complexity
Evaluating the acquisition function of ITL in round $n$ requires computing for each $\boldsymbol{x}\in\mathcal{S}$,
$\mathrm{I}(\boldsymbol{f}_{\!\!\mathcal{A}};y_{\boldsymbol{x}}\mid\mathcal{D}_{n})$
$=\frac{1}{2}\log\left(\frac{\left|\mathrm{Var}{}\left[\boldsymbol{f}_{\!\!\mathcal{A}}\mid\mathcal{D}_{n}\right]\right|}{\left|\mathrm{Var}{}\left[\boldsymbol{f}_{\!\!\mathcal{A}}\mid y_{\boldsymbol{x}},\mathcal{D}_{n}\right]\right|}\right)$ (forward)
[…]
Let $|\mathcal{S}|=m$ and $|\mathcal{A}|=k$. Then, the forward method has complexity $\mathcal{O}\left(m\cdot k^{3}\right)$. For the backward method, observe that the variances are scalar and the covariance matrix $\mathrm{Var}{}\left[\boldsymbol{f}_{\!\!\mathcal{A}}\mid\mathcal{D}_{n}\right]$ only has to be inverted once for all points $\boldsymbol{x}$. Thus, the backward method has complexity $\mathcal{O}\left(k^{3}+m\right)$.
When the size $m$ of $\mathcal{S}$ is relatively small (and hence, all points in $\mathcal{S}$ can be considered during each iteration of the algorithm), GP inference corresponds simply to computing conditional distributions of a multivariate Gaussian. The performance can therefore be improved by keeping track of the full posterior distribution over $\boldsymbol{f}_{\!\!\mathcal{S}}$ of size $\mathcal{O}\left(m^{2}\right)$ and conditioning on the latest observation during each iteration of the algorithm. In this case, after each observation the posterior can be updated at a cost of $\mathcal{O}\left(m^{2}\right)$ which does not grow with the time $n$, unlike classical GP inference.
Overall, when $m$ is small, the computational complexity of ITL is $\mathcal{O}\left(k^{3}+m^{2}\right)$. When $m$ is large (or possibly infinite) and a subset of $\tilde{m}$ points is considered in a given iteration, the computational complexity of ITL is $\mathcal{O}\left(k^{3}+\tilde{m}\cdot n^{3}\right)$, neglecting the complexity of selecting the $\tilde{m}$ candidate points. In the latter case, the computational cost of ITL is dominated by the cost of GP inference.
Khanna et al. (2017) discuss distributed and stochastic approximations of greedy algorithms to (weakly) submodular problems that are also applicable to ITL.
Appendix G Additional NN Experiments & Details
We outline the few-shot training of NNs in Algorithm 1.
Algorithm 1 Few-shot training of NNs
  Given: initialized or pre-trained model $f$, small sample $A\sim\mathcal{P}_{\!\!\mathcal{A}}$
  initialize dataset $\mathcal{D}=\emptyset$
  repeat
      sample $S\sim\mathcal{P}_{\!\mathcal{S}}$
      subsample target space $A^{\prime}\operatorname*{\smash{\overset{\mathrm{u.a.r.}}{\sim}}}A$
      initialize batch $B=\emptyset$
      compute embedding matrix $\widetilde{\boldsymbol{\Phi}}$ over domain $[S,A^{\prime}]$
      repeat $b$ times
         compute acquisition function w.r.t. $A^{\prime}$, based on $\widetilde{\boldsymbol{\Phi}}$
         add maximizer $\boldsymbol{x}\in S$ of acquisition function to $B$
         update conditional embedding matrix $\widetilde{\boldsymbol{\Phi}}$
      obtain labels for $B$ and add to dataset $\mathcal{D}$
      update $f$ using data $\mathcal{D}$
In Section G.1, we detail metrics and hyperparameters. We describe in Sections G.2 and G.3 how to compute the (initial) conditional embedding matrix $\widetilde{\boldsymbol{\Phi}}$, and in Section G.4 how to update this matrix $\widetilde{\boldsymbol{\Phi}}$ to obtain conditional embeddings for batch selection.
In Section G.5, we show that ITL and CTL significantly outperform a wide selection of commonly used heuristics. In Sections G.6 and G.7, we conduct additional experiments and ablations.
G.1 Experiment Details
Table 1: Hyperparameter summary of NN experiments.
MNIST CIFAR-100
$\rho$ $0.01$ $1$
$M$ $30$ $100$
$m$ $3$ $10$
$k$ 1 00010001\,0001 000 1 00010001\,0001 000
batch size $b$ $1$ $10$
# of epochs (*) [^15] We train until convergence with respect to oracle validation accuracy. $5$
learning rate $0.001$ $0.001$
We evaluate the accuracy with respect to $\mathcal{P}_{\!\!\mathcal{A}}$ using a Monte Carlo approximation with out-of-sample data:
$\text{accuracy}(\boldsymbol{\widehat{\theta}})\approx\mathbb{E}_{(\boldsymbol{x},y)\sim\mathcal{P}_{\!\!\mathcal{A}}}{}\mathbbm{1}\{{y=\operatorname*{arg\,max}_{i}f_{i}(\boldsymbol{x};\boldsymbol{\widehat{\theta}})}\}.$
We provide an overview of the hyperparameters used in our NN experiments in Table 1. The effect of noise standard deviation $\rho$ is small for all tested $\rho\in[1,100]$ (cf. ablation study in Table 2). [^16] We use a larger noise standard deviation $\rho$ in CIFAR-100 to stabilize the numerics of batch selection via conditional embeddings (cf. Table 2). $M$ denotes the size of the sample ${A\sim\mathcal{P}_{\!\!\mathcal{A}}}$. In each iteration, we select the target space $\mathcal{A}\leftarrow A^{\prime}$ as a random subset of $m$ points from $A$ [^17] This appears to improve the training, likely because it prevents overfitting to peculiarities in the finite sample $A$. We observe that choosing $m$ to match the number of classes yields good performance.
During each iteration, we select the batch $B$ by choosing the top-$b$ points according to the decision rule from a random sample from $\mathcal{P}_{\!\mathcal{S}}$ of size $k$ [^18] In large-scale problems, the work of Coleman et al. (2022) suggests to use an (approximate) nearest neighbor search to select the (large) candidate set rather than sampling u.a.r. from $\mathcal{P}_{\!\mathcal{S}}$. This can be a viable alternative to simply increasing $k$ and suggests future work.
Since we train the MNIST model from scratch, we train from random initialization until convergence on oracle validation accuracy. [^19] That is, to stop training as soon as accuracy on a validation set from $\mathcal{P}_{\!\!\mathcal{A}}$ decreases in an epoch. We do this to stabilize the learning curves, and provide the least biased (due to the training algorithm) results. For CIFAR-100, we train for $5$ epochs (starting from the previous iterations’ model) which we found to be sufficient to obtain good performance.
We use the ADAM optimizer (Kingma & Ba, 2014). In our CIFAR-100 experiments, we use a pre-trained EfficientNet-B0 (Tan & Le, 2019) model, and fine-tune the final and penultimate layers. We freeze earlier layers to prevent overfitting to the few-shot training data.
To prevent numerical inaccuracies when computing the ITL objective, we optimize
$\mathrm{I}(\boldsymbol{y}_{\!\mathcal{A}};y_{\boldsymbol{x}}\mid\mathcal{D}_{n-1})=\frac{1}{2}\log\left(\frac{\mathrm{Var}{}\!\left[y_{\boldsymbol{x}}\ \middle|\ \mathcal{D}_{n-1}\right]}{\mathrm{Var}{}\!\left[y_{\boldsymbol{x}}\ \middle|\ \boldsymbol{y}_{\!\mathcal{A}},\mathcal{D}_{n-1}\right]}\right)$ (24)
instead of Equation 1, which amounts to adding $\rho^{2}$ to the diagonal of the covariance matrix before inversion. This appears to improve numerical stability, especially when using gradient embeddings. [^20] In our experiments, we observe that the effect various choices of $\rho$ on this slight adaptation of the ITL decision rule has negligible impact on performance. The more prominent effect of $\rho$ appears to arise from the batch selection via conditional embeddings (cf. Table 2).
G.2 Gradient vs. Latent Embeddings
[…]
[]nns_embeddings
Figure 4: Comparison of gradient (“G-”) and latent embeddings (“L-”).
Using a neural network to parameterize $f$, we evaluate the following canonical approximations of $f$ by a stochastic process:
  1. 1.
    Latent Embeddings: Taking the latent representation $\boldsymbol{\phi}(\boldsymbol{x})$ from the penultimate layer and collecting them as rows in the design matrix $\boldsymbol{\Phi}$ of inputs $X$, we can view the last layer as the linear function ${\boldsymbol{f}_{\!\!X}=\boldsymbol{\Phi}\boldsymbol{\theta}}$. Approximating the weights by ${\boldsymbol{\theta}\sim\mathcal{N}({}\boldsymbol{\mu},\boldsymbol{\Sigma})}$ implies $\boldsymbol{f}_{\!\!X}\sim\mathcal{N}({}\boldsymbol{\Phi}\boldsymbol{\mu},\boldsymbol{\Phi}\boldsymbol{\Sigma}\boldsymbol{\Phi}^{\top})$. Interpreting the early layers as a feature encoder, thus uses the low-dimensional feature map akin to random feature methods (Rahimi & Recht, 2007).
  2. 2.
    Gradient Embeddings: Alternatively, one can approximate ${\boldsymbol{f}_{\!\!X}\sim\mathcal{N}({}\boldsymbol{\mu},\boldsymbol{J}\boldsymbol{\Sigma}\boldsymbol{J}^{\top})}$ where $\boldsymbol{J}=\boldsymbol{\nabla}_{\!\!\boldsymbol{\theta}}\,\boldsymbol{f}(X;\boldsymbol{\theta})$ is the Jacobian of NN parameters $\boldsymbol{\theta}$, effectively taking the gradient embeddings $\boldsymbol{g}(\boldsymbol{x})=\boldsymbol{\nabla}_{\!\!\boldsymbol{\theta}}\,f(\boldsymbol{x};\boldsymbol{\theta})$ as feature representations (Khan et al., 2019; Holzmüller et al., 2023). This kernel is called the (empirical) neural tangent kernel (NTK) and the posterior mean of this GP approximates wide NNs trained with gradient descent (Arora et al., 2019). Kassraie & Krause (2022) derive bounds of $\gamma_{n}$ under this kernel.
    Rather than the NTK, we use in our experiments the gradient embeddings
    $\boldsymbol{g}(\boldsymbol{x})=\left.\boldsymbol{\nabla}_{\!\!\boldsymbol{\theta^{\prime}}}\,\ell_{\mathrm{CE}}(f(\boldsymbol{x};\boldsymbol{\theta}),\hat{y}(\boldsymbol{x}))\right\rvert_{\boldsymbol{\theta}=\boldsymbol{\widehat{\theta}}}$
    where $\boldsymbol{\theta^{\prime}}$ refers to the parameters of the final output layer, $\boldsymbol{\widehat{\theta}}$ are the current parameter estimates, $\smash{\hat{y}(\boldsymbol{x})=\operatorname*{arg\,max}_{i}f_{i}(\boldsymbol{x};\boldsymbol{\widehat{\theta}})}$ are the associated predicted labels, and $\ell_{\mathrm{CE}}$ denotes the cross-entropy loss. This gradient embedding captures the potential update direction upon observing a new point (Ash et al., 2020). Moreover, Ash et al. (2020) show that for most neural networks, the norm of these gradient embeddings are a conservative lower bound to the norm assumed by taking any other proxy label $\hat{y}(\boldsymbol{x})$.
In the following, we denote the design matrix of inputs $X$ w.l.o.g. by $\boldsymbol{\Phi}$ (collecting the embeddings of all $\boldsymbol{x}\in X$ as rows), and the covariance matrix of $\boldsymbol{f}_{\!\!X}$ by $\widetilde{\boldsymbol{\Phi}}\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\boldsymbol{\Phi}\boldsymbol{\Sigma}\boldsymbol{\Phi}^{\top}$.
Unless otherwise specified, we use gradient embeddings in our experiments. We compare latent and gradient embeddings in Figure 4 and observe only negligible differences in performance.
G.3 Towards Quantifying Epistemic Uncertainty
[…]
[]nns_uncertainty_quantification
Figure 5: Comparison uncertainty quantification (i.e., estimation of $\boldsymbol{\Sigma}$) via a Laplace approximation (LA, Daxberger et al. (2021)) using a Kronecker factored log-likelihood Hessian approximation (Martens & Grosse, 2015). The results are shown for the MNIST experiment. We do not observe a performance improvement beyond the trivial approximation $\boldsymbol{\Sigma}=\boldsymbol{I}$.
A straightforward and common approximation of the uncertainty about NN weights is given by $\boldsymbol{\Sigma}=\boldsymbol{I}$, and we use this approximation throughout our experiments.
The poor performance of UnSa (cf. Section G.5) with this approximation suggests that with more sophisticated models, the performance of ITL and CTL can be further improved. Various approximations of $\boldsymbol{\Sigma}$ based on Laplace approximation (Daxberger et al., 2021), variational inference (Blundell et al., 2015), or Markov chain Monte Carlo (Maddox et al., 2019) have been studied. We also evaluate Laplace approximation (LA, Daxberger et al. (2021)) for estimating $\boldsymbol{\Sigma}$ but see no improvement (cf. Figure 5). Nevertheless, we believe that uncertainty quantification is a promising direction for future work, with the potential to improve performance of ITL and its variations substantially.
G.4 Batch Selection via Conditional Embeddings
We will refer to Equation 5 as BaCE, short for batch selection via conditional embeddings. BaCE can be implemented efficiently using the Gaussian approximation of $\boldsymbol{f}_{\!\!X}$ from Section G.2 by iteratively conditioning on the previously selected points $\boldsymbol{x}_{n,1:i-1}$, and updating the embedding matrix $\smash{\widetilde{\boldsymbol{\Phi}}}$ using the closed-form formula for the variance of conditional Gaussians:
$\widetilde{\boldsymbol{\Phi}}\leftarrow\widetilde{\boldsymbol{\Phi}}-\frac{1}{\widetilde{\boldsymbol{\Phi}}_{j,j}+\rho^{2}}\widetilde{\boldsymbol{\Phi}}_{X,j}\widetilde{\boldsymbol{\Phi}}_{j,X}$ (25)
where $j$ denotes the index of the selected $\boldsymbol{x}_{n,i}$ within $X$ and $\rho^{2}$ is the noise variance. Note that $\smash{\widetilde{\boldsymbol{\Phi}}}_{j,j}$ is a scalar and $\smash{\widetilde{\boldsymbol{\Phi}}}_{X,j}$ is a row vector, and hence, this iterative update can be implemented efficiently.
We remark that Equations 4 and 5 are natural extensions of previous non-adaptive active learning methods, which typically maximize some notion of “distance” between points in the batch, to the “directed” setting (Ash et al., 2020; Zanette et al., 2021; Holzmüller et al., 2023; Pacchiano et al., 2024). BaCE simultaneously maximizes “distance” between points in a batch and minimizes “distance” to points in $\mathcal{A}$.
G.5 Baselines
[…]
[]nns_undirected
Figure 6: Comparison of baselines for the experiment of Figure 2.
Denote the softmax distribution over labels $i$ at inputs $\boldsymbol{x}$ by
$p_{i}(\boldsymbol{x};\boldsymbol{\widehat{\theta}})\propto\exp(f_{i}(\boldsymbol{x};\boldsymbol{\widehat{\theta}})).$
We compare to the following commonly used heuristic “undirected” decision rules:
  • BADGE (Ash et al., 2020) which selects a batch according to the kmeans++ initialization scheme, using the gradient embeddings $\boldsymbol{g}(\boldsymbol{x})$ as features.
  • MaxMargin (Settles & Craven, 2008):
    $\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{S}}p_{1}(\boldsymbol{x};\boldsymbol{\widehat{\theta}}_{n-1})-p_{2}(\boldsymbol{x};\boldsymbol{\widehat{\theta}}_{n-1})$
    where $p_{1}$ and $p_{2}$ are the two largest class probabilities.
  • MaxEntropy (Settles & Craven, 2008):
    $\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{S}}\mathrm{H}[p(\boldsymbol{x};\boldsymbol{\widehat{\theta}}_{n-1})].$
  • Uncertainty sampling (UnSa) which with embeddings $\boldsymbol{\phi}_{n-1}$ corresponds to:
    $\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{S}}\sigma_{n-1}^{2}(\boldsymbol{x})=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{S}}\left\|\boldsymbol{\phi}_{n-1}(\boldsymbol{x})\right\|_{2}^{2}.$
    We evaluate UnSa with gradient embeddings. This acquisition function is similar to previously used “embedding length” or “gradient length” heuristics (Settles & Craven, 2008).
We summarize the results in Figure 6. We observe that most “undirected” decision rules perform worse (and often significantly so) than Random. This is likely due to frequently selecting points from the support of $\mathcal{P}_{\!\mathcal{S}}$ which are not in the support of $\mathcal{P}_{\!\!\mathcal{A}}$ since the points are “adversarial examples” that the model $\boldsymbol{\widehat{\theta}}$ is not trained to perform well on. In the case of MNIST, this can also partially be attributed to the well-known “cold-start problem” (Gao et al., 2020).
We also compare to the following “directed” decision rules:
  • CosineSimilarity (Settles & Craven, 2008) selects $\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{S}}\angle_{\boldsymbol{\phi}_{n-1}}(\boldsymbol{x},\mathcal{A})$ where
    $\angle_{\boldsymbol{\phi}}(\boldsymbol{x},\mathcal{A})\operatorname*{\smash{\overset{\mathrm{def}}{=}}}\sum_{\boldsymbol{x^{\prime}}\in\mathcal{A}}\frac{\boldsymbol{\phi}(\boldsymbol{x})^{\top}\boldsymbol{\phi}(\boldsymbol{x^{\prime}})}{\left\|\boldsymbol{\phi}(\boldsymbol{x})\right\|_{2}\left\|\boldsymbol{\phi}(\boldsymbol{x^{\prime}})\right\|_{2}}.$
  • InformationDensity (Settles & Craven, 2008) is defined as the multiplicative combination of MaxEntropy and CosineSimilarity:
    $\boldsymbol{x}_{n}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{S}}\mathrm{H}[p(\boldsymbol{x};\boldsymbol{\widehat{\theta}}_{n-1})]\cdot\left(\frac{1}{|\mathcal{A}|}\angle_{\boldsymbol{\phi}_{n-1}}(\boldsymbol{x},\mathcal{A})\right)^{\beta}$
    where $\beta>0$ controls the relative importance of both terms. We set $\beta=1$ in our experiments.
We evaluate both baselines using latent embeddings (cf. Section G.2), as is typical.
G.6 Additional experiments
[…]
[]nns_imbalanced_train
Figure 7: Imbalanced $\mathcal{P}_{\!\mathcal{S}}$ experiment. This amounts to making the signal in $\mathcal{P}_{\!\mathcal{S}}$ about $\mathcal{P}_{\!\!\mathcal{A}}$ more sparse, simulating more difficult large-scale learning tasks.
[…]
[]nns_imbalanced_test
Figure 8: Imbalanced $A\sim\mathcal{P}_{\!\!\mathcal{A}}$ experiment.
We conduct two additional experiments:
  1. 1.
    Imbalanced $\mathcal{P}_{\!\mathcal{S}}$ (Figure 7): We artificially remove $80\%$ of the support of $\mathcal{P}_{\!\!\mathcal{A}}$ from $\mathcal{P}_{\!\mathcal{S}}$. For example, in case of MNIST, we remove $80\%$ of the images with labels $3$, $6$, and $9$ from $\mathcal{P}_{\!\mathcal{S}}$. This makes the learning task more difficult, as $\mathcal{P}_{\!\!\mathcal{A}}$ is less represented in $\mathcal{P}_{\!\mathcal{S}}$, meaning that the “targets” are more sparse. The trend of ITL outperforming CTL which outperforms Random is even more pronounced in this setting.
  2. 2.
    Imbalanced $A\sim\mathcal{P}_{\!\!\mathcal{A}}$ (Figure 8): We artificially remove $50\%$ of part of the support of $\mathcal{P}_{\!\!\mathcal{A}}$ while generating $A\sim\mathcal{P}_{\!\!\mathcal{A}}$ to evaluate the robustness of ITL and CTL in presence of an imbalanced target space $\mathcal{A}$. Concretely, in case of MNIST, we remove $50\%$ of the images with labels $3$ and $6$ from $A$. In case of CIFAR-100, we remove $50\%$ of the images with labels $\{0,\dots,4\}$ from $A$. We still observe the same trends as in the other experiments.
G.7 Ablation study of noise standard deviation $\rho$
In Table 2, we evaluate the CIFAR-100 experiment with different noise standard deviations $\rho$. We observe that the performance of batch selection via conditional embeddings drops (mostly for the less numerically stable gradient embeddings) if $\rho$ is too small, since this leads to numerical inaccuracies when computing the conditional embeddings. Apart from this, the effect of $\rho$ is negligible.
Table 2: Ablation study of noise standard deviation $\rho$ in the CIFAR-100 experiment. We list the accuracy after $100$ rounds per decision rule, with its standard error over $10$ random seeds. “(top-$b$)” denotes variants where batches are selected by taking the top-$b$ points according to the decision rule rather than using batch selection via conditional embeddings. Shown in bold are the best performing decision rules, and shown in italics are results due to numerical instability.
$\rho$ $0.0001$ $0.01$ $1$ $100$
G-ITL $\mathit{78.26\pm 1.40}$ $\mathit{79.12\pm 1.19}$ $\mathbf{87.16\pm 0.29}$ $\mathbf{87.18\pm 0.28}$
L-ITL $\mathbf{87.52\pm 0.48}$ $\mathbf{87.52\pm 0.41}$ $\mathbf{87.53\pm 0.35}$ $86.47\pm 0.22$
G-CTL $\mathit{58.68\pm 2.11}$ $\mathit{81.44\pm 1.04}$ $86.52\pm 0.44$ $\mathbf{86.92\pm 0.56}$
L-CTL $\mathbf{86.40\pm 0.71}$ $\mathbf{86.38\pm 0.75}$ $86.00\pm 0.69$ $84.78\pm 0.39$
G-ITL (top-$b$) $85.84\pm 0.54$ $85.92\pm 0.52$ $85.84\pm 0.54$ $85.55\pm 0.46$
L-ITL (top-$b$) $85.44\pm 0.58$ $85.46\pm 0.54$ $85.44\pm 0.59$ $85.29\pm 0.36$
G-CTL (top-$b$) $82.27\pm 0.67$ $82.27\pm 0.67$ $82.27\pm 0.67$ $82.27\pm 0.67$
L-CTL (top-$b$) $80.73\pm 0.68$ $80.73\pm 0.68$ $80.73\pm 0.68$ $80.73\pm 0.68$
BADGE $83.24\pm 0.60$ $83.24\pm 0.60$ $83.24\pm 0.60$ $83.24\pm 0.60$
InformationDensity $79.24\pm 0.51$ $79.24\pm 0.51$ $79.24\pm 0.51$ $79.24\pm 0.51$
Random $82.49\pm 0.66$ $82.49\pm 0.66$ $82.49\pm 0.66$ $82.49\pm 0.66$
Table 3: Magnitudes of $\gamma_{n}$ for common kernels. The magnitudes hold under the assumption that $\mathcal{X}$ is compact. Here, $B_{\nu}$ is the modified Bessel function. We take the magnitudes from Theorem 5 of Srinivas et al. (2009) and Remark 2 of Vakili et al. (2021). The notation $\widetilde{\mathcal{O}}\left(\cdot\right)$ subsumes log-factors. For $\nu=1/2$, the Matérn kernel is equivalent to the Laplace kernel. For $\nu\to\infty$, the Matérn kernel is equivalent to the Gaussian kernel. The functions sampled from a Matérn kernel are $\lceil\nu\rceil-1$ mean square differentiable.
Kernel $k(\boldsymbol{x},\boldsymbol{x^{\prime}})$ $\gamma_{n}$
Linear $\boldsymbol{x}^{\top}\boldsymbol{x^{\prime}}$ $\widetilde{\mathcal{O}}\left(d\log(n)\right)$
Gaussian $\exp\left(-\frac{\left\|\boldsymbol{x}-\boldsymbol{x^{\prime}}\right\|_{2}^{2}}{2h^{2}}\right)$ $\widetilde{\mathcal{O}}\left(\log^{d+1}(n)\right)$
Laplace $\exp\left(-\frac{\left\|\boldsymbol{x}-\boldsymbol{x^{\prime}}\right\|_{1}}{h}\right)$ $\widetilde{\mathcal{O}}\left(n^{\frac{d}{1+d}}\log^{\frac{1}{1+d}}(n)\right)$
Matérn $\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}\left\|\boldsymbol{x}-\boldsymbol{x^{\prime}}\right\|_{2}}{h}\right)^{\nu}B_{\nu}\left(\frac{\sqrt{2\nu}\left\|\boldsymbol{x}-\boldsymbol{x^{\prime}}\right\|_{2}}{h}\right)$ $\widetilde{\mathcal{O}}\left(n^{\frac{d}{2\nu+d}}\log^{\frac{2\nu}{2\nu+d}}(n)\right)$